### v1

# Optimal control under unknown intensity with Bayesian learning

Abstract
We consider an optimal control problem inspired by neuroscience, where the dynamics is driven by a Poisson process with a controlled stochastic intensity and an uncertain parameter. Given a prior distribution for the unknown parameter, we describe its evolution according to Bayes’ rule. We reformulate the optimization problem using Girsanov’s theorem and establish a dynamic programming principle. Finally, we characterize the value function as the unique viscosity solution to a finite-dimensional Hamilton-Jacobi-Bellman equation, which can be solved numerically.
Keywords: Optimal control, Bayesian filtering, Point processes.
Mathematics Subject Classification 62M20, 49L20, 49L25.
1 Introduction
Let $(Y_{t})$ represent the membrane potential of a spiking neuron. The neuron emits “spikes” randomly at rate $f_{\Lambda}(Y_{t})$. When a spike occurs, the potential is reset to zero. Finally, between the jumps, the dynamics is given by the ODE
$\dot{Y_{t}}=b(Y_{t})+\gamma_{t}.$
In this model, the functions $(\lambda,y)\mapsto f_{\lambda}(y)\in\mathbb{R}_{+}$ and $y\mapsto b(y)\in\mathbb{R}$ are deterministic and known. In addition, $\Lambda$ is a random variable with known distribution $\mu$, supported on $\mathbb{R}_{+}$, but the realization of $\Lambda$ is unknown. Our objective is to estimate as accurately as possible the value of $\Lambda$. In addition, the input current $(\gamma_{t})$ is a control and can be chosen in order to improve the estimation. Crucially, we only observe the jumping times of $(Y_{t})$ and therefore $(\gamma_{t})$ is only allowed to causally depend on the spiking times up to time $t$. The question we answer in this article can be summarized as follows:
How to choose the stochastic control $(\gamma_{t})$ to optimally estimate the unknown parameter?
This leads us to consider an optimal control problem driven by a counting process with unknown intensity. Our approach is applicable beyond neuroscience, where optimal control over processes with unknown parameters is required, such as in online learning scenarios. We present a mathematical framework to tackle this class of questions.
Such problems have been widely studied in the discrete-time stochastic optimal control literature (see, e.g., [6, 7]). A key approach is to assign a prior distribution to the unknown parameter, which is updated at each time step using Bayes’ rule. In the continuous-time setting, particularly within the impulse control framework, these issues have been addressed by [2] in a Brownian framework and extended by [1] in a Poisson framework. As in the discrete-time setting, a Bayesian framework is used: the unknown parameter is sampled from a prior distribution and a stochastic process in the space of probability measures is introduced, known as the posterior distribution. This approach leads to a dynamic programming principle and a characterization of the value function via a quasi-variational parabolic equation interpreted in the sense of viscosity solutions.
In [8] (lecture given at Collège de France) and [4] (with an application in asset management), the authors consider instances of the following diffusive case:
$dX_{t}=\Lambda b(X_{t},\gamma_{t})\mathrm{d}t+\sigma(X_{t})\mathrm{d}W_{t}.$
In this model, one observes the process $(X_{t})$ and the parameter $\Lambda$ is unknown. A key idea introduced in these works is to reformulate the problem using Girsanov’s theorem. This permits removing the drift and to make explicit the dependence of the problem on the unknown parameter via the exponential martingale involved in Girsanov’s transformation. In addition, in this linear case in $\Lambda$ or for specific classes of prior distribution known as conjugate families, they derive a finite dimensional HJB equation satisfied by the value function.
In the setting we introduce, a new mathematical challenge arises because the conditional distribution of the unknown parameters evolves continuously and also exhibits jumps: the jumps of the conditional distribution occurs precisely at the jumping times of the underlying point process; with intensity depending on the unknown parameter. Another new important difficulty appear: when the control $(\gamma_{t})$ is modified, the jumps of the process are also modified. As the controller is only allowed to observe the spiking times, it is not clear to know at first sight what are the admissible controls. In particular, it is not possible to use the same control $(\gamma_{t})$ for two different initial conditions $y$ and $y^{\prime}$ of $(Y_{t})$ because changing the initial condition might change the spiking times and therefore a control which was valid for $y$ is not valid anymore for $y^{\prime}$. This feature makes particularly difficult to obtain apriori regularity estimates on the value function. As in [8, 4], Girsanov’s theorem plays an important role in our work in addressing these difficulties. Another difficulty is due to the fact that the measure space is infinite-dimensional, and the posterior distribution evolves continuously over time on this space. In the case where the intensity of the original point process is linearly dependent on the unknown parameter, we show that the posterior distribution is confined to a two-dimensional subspace of $\mathbb{R}^{2}$. We further show that the value function associated with the optimal control problem is the unique solution to a specific finite-dimensional Hamilton-Jacobi-Bellman (HJB) equation in the viscosity sense.
We now describe briefly our main contributions and the organization of this paper.
The control problem is introduced in Section 2.1, where we also characterize the admissible controls and show that the optimization problem is independent of the probability space we started with. In Section 2.2, we reformulate the problem using Girsanov’s theorem for point processes with stochastic intensity and show the equivalence between the original and the new formulation. In Section 2.3, the value function is introduced, and in Section 2.4, we show how to obtain a finite-dimensional representation of the value function provided that function $f_{\lambda}(y)$ is linear in $\lambda$, that is $f_{\lambda}(y)=\lambda g(y)$. This is analogous to the finite-dimensional reduction obtained in [8, 4] in the diffusive setting with a linear drift. Thereafter, we concentrate our study on this finite-dimensional problem.
In Section 3.1, we prove that the value function is Lipschitz continuous with respect to the space variables. To proceed, we study in detail the exponential martingale involved in the Girsanov’s transformation. Using these apriori regularity estimates, we prove in Section 3.2 a first dynamic programming principle with deterministic stopping times. Leveraging this result, we show in Section 3.3 that the value function is continuous with respect to time. Finally, in Section 3.4 we provide a general dynamic programming principle that allows for general stopping times.
In Section 4, we show that the value function is the unique viscosity solution of a certain HJB equation, see equation (4.1). This first order HJB equation has zero order terms which are inherent to the jumps in the underlying dynamics. We rely on the methodology of [9] to handle these terms. A specific difficulty appears in the proof of our comparison result, due to the fact that the controls are not bounded, neither are the space variables.
Finally, in Section 5, we present several numerical illustrations and explain the behavior of the optimal control.
2 The control problem
2.1 The optimization problem
Let $\mu\in{\mathcal{P}}_{2}(\mathbb{R}_{+})$ be a probability measure supported on $\mathbb{R}_{+}$ with a finite second moment. Consider functions $b$ and $f$ such that:
Assumption 2.1.
The function $b:\mathbb{R}\rightarrow\mathbb{R}$ is Lipschitz and $f:\mathbb{R}_{+}\times\mathbb{R}\rightarrow\mathbb{R}_{+}$ is continuous and bounded for each $\lambda\geq 0$:
$\sup_{y\in\mathbb{R}}f_{\lambda}(y)<\infty.$
Let $(\Omega,{\mathcal{F}},({\mathcal{F}}_{t}),\mathbb{P})$ be a filtered probability space satisfying the usual conditions. Fix a time horizon $T>0$. Consider $(\gamma_{t})_{t\in[0,T]}$ an $({\mathcal{F}}_{t})$-progressively measurable process such that
$\mathbb{P}(\mathrm{d}\omega)a.s,\quad\gamma\in L^{2}([0,T]).$
Definition 2.2.
We consider a triple $(\Lambda,(N_{t}),(Y_{t}))$ which is a solution of:
  1. 1.
    $\Lambda$ is an ${\mathcal{F}}_{0}$-measurable random variable with probability distribution $\mu$.
  2. 2.
    $(N_{t})$ is an $({\mathcal{F}}_{t})$ point process with stochastic intensity $f_{\Lambda}(Y_{t-})$.
  3. 3.
    $(Y_{t})$ solves:
    $Y_{t}=y+\int_{0}^{t}(b(Y_{u})+\gamma_{u})\mathrm{d}u-\int_{0}^{t}Y_{u-}\mathrm{d}N_{u},\quad t\in[0,T].$
We assume that the probability space is large enough so that for each $\gamma$ as above, there exists at least one triple $(\Lambda,(N_{t}),(Y_{t}))$ defined on this probability space, satisfying the three conditions above.
Let ${\mathcal{F}}^{N}_{t}$ denote the canonical filtration of $(N_{t})_{t\geq 0}$:
${\mathcal{F}}^{N}_{t}=\sigma\{N_{s},~{}s\leq t\}.$
Let $(\tau_{k})_{k\geq 1}$ denote the successive jump times of $(N_{t})_{t\geq 0}$. We set by convention that $\tau_{0}=0$.
Definition 2.3.
A control $\gamma$ is said to be admissible if it is predictable with respect to the filtration $({\mathcal{F}}^{N}_{t})_{t\geq 0}$.
We denote by ${\mathcal{A}}$ the set of admissible controls. Furthermore, we consider the posterior distribution of $\Lambda$ given the observed spiking times, denoted by:
$M^{\gamma}_{t}:={\mathcal{L}}(\Lambda~{}|~{}{\mathcal{F}}^{N}_{t}).$
Let $\kappa>0$. Our optimization problem is formulated as follows:
$\inf_{\gamma\in{\mathcal{A}}}\mathbb{E}\left[\frac{1}{2}\int_{0}^{T}{\gamma^{2}_{s}\mathrm{d}s}+\kappa\text{Var}(M^{\gamma}_{T})\right].$ (2.1)
In other words, our goal is to minimize the variance of the posterior distribution; the quadratic cost in the control is a regularization to prevent for too large controls. We now rewrite this optimization problem in a more explicit way, specifically to demonstrate that the value defined by (2.1) is independent of the probability space $(\Omega,{\mathcal{F}},({\mathcal{F}}_{t}),\mathbb{P})$. The following proposition characterizes the admissible controls.
Proposition 2.4.
Let $\gamma$ be a ${\mathcal{F}}^{N}$-predictable process such that $\mathbb{P}(\mathrm{d}\omega)$ a.s., $\gamma\in L^{2}([0,T])$. There exists a sequence of random variables:
$\Gamma_{k}:\Omega\rightarrow L^{2}([0,T])$
such that $\Gamma_{k}$ is ${\mathcal{F}}^{N}_{\tau_{k}}$-measurable and
$\mathbb{P}(\mathrm{d}\omega)\times\mathrm{d}t~{}a.s.,\quad\gamma_{t}(\omega)=\sum_{k\geq 0}\Gamma_{k}(t)\mathbbm{1}_{(\tau_{k},\tau_{k+1}]}(t).$
Proof.
See [5, A2, Th. 34]. ∎
Consider $\mathbb{D}([0,T])$, the Skorokhod space of càdlàg functions on $[0,T]$. Since each random variable $\Gamma_{k}$ is ${\mathcal{F}}^{N}_{\tau_{k}}$-measurable, there exists a deterministic and measurable function $\Psi_{k}:\mathbb{D}([0,T])\rightarrow L^{2}([0,T])$ such that:
$\Gamma_{k}=\Psi_{k}(N_{\cdot\wedge\tau_{k}}).$
This establishes a one-to-one correspondence between an admissible control $\gamma$ and a collection of measurable functions $(\Psi_{k})_{k\geq 0}$. Write =$\mathbb{N}=\{0,1,\cdots\}$ for the set of natural integers. We denote by $\tilde{A}$ the set of all the measurable functions from $D([0,T])$ to $(L^{2}([0,T]))^{\mathbb{N}}$.
In addition, there exists a deterministic function $V:D([0,T])\rightarrow\mathbb{R}$ such that
$\mathbb{P}(\mathrm{d}\omega)~{}a.s.,\quad\text{Var}(M^{\gamma}_{T})=\mathbb{E}[\Lambda^{2}~{}|~{}{\mathcal{F}}^{N}_{T}]-\mathbb{E}^{2}[\Lambda~{}|~{}{\mathcal{F}}^{N}_{T}]=V(N).$
Remark 2.5.
We will see in Corollary 2.9 below that this function $V$ is explicit.
Given $\Psi\in\tilde{A}$, consider $(\Lambda,N_{t},V_{t})$ which satisfies the three properties outlined in Definition 2.2, with the control given by $\gamma=\sum_{k\geq 0}\Psi_{k}(N_{\cdot\wedge\tau_{k}})\mathbbm{1}_{(\tau_{k},\tau_{k+1}]}$. These three conditions fully determine the law of $(N_{t})_{t\in[0,T]}$. We denote by $Q_{\Psi,\mu}$ this probability law:
$Q_{\Psi,\mu}={\mathcal{L}}((N_{t})_{t\in[0,T]}).$
Thus, our optimization problem (2.1) is equal to:
$\inf_{\Psi\in\tilde{A}}\mathbb{E}\left[\sum_{k\geq 0}\frac{1}{2}\int_{0}^{T}{\Psi^{2}_{k}(N_{\cdot\wedge\tau_{k}})\mathbbm{1}_{(\tau_{k},\tau_{k+1}]}(s)\mathrm{d}s}+\kappa V(N))\right]$
$=\inf_{\Psi\in\tilde{A}}\int\left[\sum_{k\geq 0}\frac{1}{2}\int_{0}^{T}{\Psi^{2}_{k}(\tilde{n}_{\cdot\wedge\tilde{\tau}_{k}})\mathbbm{1}_{(\tilde{\tau}_{k},\tilde{\tau}_{k+1}]}(s)\mathrm{d}s}+\kappa V(\tilde{n}))\right]Q_{\Psi,\mu}(\mathrm{d}\tilde{n}).$
In the last expression, the $\tilde{\tau}_{k}$ denotes the jumps of $\tilde{n}$. The key observation here is that this expression is entirely independent of the original probability space from which we began.
2.2 Girsanov’s theorem and the posterior distribution
We observed that our optimization problem does not depend on the initial probability space. We propose here a specific and convenient construction of these objects, using Girsanov’s theorem.
We consider $(\Omega,{\mathcal{F}},({\mathcal{F}}_{t}),\mathbb{Q})$ a probability space such that $(N_{t})$ is a standard $({\mathcal{F}}_{t})$-Poisson process with rate $1$. We denote by $(\tau_{k})$ the successive jump times of $(N_{t})$. We write ${\mathcal{F}}^{N}_{t}=\sigma(N_{r},r\leq t)$ for the natural filtration of $N$.
Let ${\mathcal{A}}$ be the collection of all the ${\mathcal{F}}^{N}$-predictable processes $\gamma$ such that
$a.s.,\quad\int_{0}^{T}\gamma^{2}_{u}\mathrm{d}u<\infty.$
Given $\gamma\in{\mathcal{A}}$ and $y\in\mathbb{R}$, we consider $(Y_{t})$ the solution of the SDE
$Y_{t}=y+\int_{0}^{t}(b(Y_{u})+\gamma_{u})\mathrm{d}u-\int_{0}^{t}Y_{u-}\mathrm{d}N_{u}.$
For all $\lambda\in\mathbb{R}_{+}$, we define:
$L_{t}(\lambda)$ $=\prod_{0<\tau_{k}\leq t}f_{\lambda}(Y_{\tau_{k}-})\exp\left(t-\int_{0}^{t}f_{\lambda}(Y_{u})\mathrm{d}u\right).$
Lemma 2.6.
Let $\tilde{N}_{t}=N_{t}-t$. The equation
$L_{t}=1+\int_{0}^{t}L_{u-}[f_{\lambda}(Y_{u-})-1]\mathrm{d}\tilde{N}_{u}$
has a unique locally bounded solution ($\sup_{u\in[0,t]}|L_{u}|<\infty~{}a.s.$) and this solution is $L_{t}(\lambda)$ as defined above. In addition, $t\mapsto L_{t}(\lambda)$ is a $(\mathbb{Q},({\mathcal{F}}^{N}_{t}))$-local martingale.
Proof.
This follows from [5][Th. A.T4] ∎
Consider now $\Lambda$ a ${\mathcal{F}}_{0}$-random variable, with probability law $\mu$. As $(N_{t})$ is a $({\mathcal{F}}_{t})$ standard Poisson process, it holds that $\Lambda$ and $(N_{t})_{t\geq 0}$ are independent under $\mathbb{Q}$.
Lemma 2.7.
Under Assumption 2.1, $\mathbb{E}_{\mathbb{Q}}L_{T}(\Lambda)=1$ and so $(L_{t}(\Lambda))$ is a $(\mathbb{Q},({\mathcal{F}}_{t}))$-martingale.
Proof.
By assumptions, for all $\lambda\in\mathbb{R}_{+}$,
$C(\lambda):=\sup_{x\geq 0}f_{\lambda}(x)<\infty.$
We first prove that $\mathbb{E}_{\mathbb{Q}}L_{T}(\lambda)=1$ for all $\lambda\in\mathbb{R}_{+}$. We have $L_{t}\leq C(\lambda)^{N_{t}}e^{t}$ and $\lvert f_{\lambda}(Y_{t-})-1\rvert\leq C(\lambda)+1$. Therefore,
$\mathbb{E}_{\mathbb{Q}}\int_{0}^{T}L_{u-}(\lambda)\lvert f_{\lambda}(Y_{u-})(\lambda)-1\rvert\mathrm{d}u\leq\int_{0}^{T}e^{sC(\lambda)}(1+C(\lambda))\mathrm{d}u<\infty.$
By [5, T8, p. 27], this implies that $\mathbb{E}_{\mathbb{Q}}L_{T}(\lambda)=1$. Finally note that there exists a measurable function $\Phi$ such that
$L_{T}(\lambda)=\Phi(N_{\cdot\wedge T},\lambda).$
On $\mathbb{Q}$, $(N_{t})_{t\geq 0}$ and $\Lambda$ are independent. Therefore,
$\mathbb{E}_{\mathbb{Q}}[\Phi(N_{\cdot\wedge T},\Lambda)~{}|~{}\Lambda]=\Psi(\Lambda),$
where $\Psi(\lambda):=\mathbb{E}_{\mathbb{Q}}\Phi(N_{\cdot\wedge T},\lambda)=\mathbb{E}_{\mathbb{Q}}L_{T}(\lambda)=1$. We deduce that $\mathbb{E}_{\mathbb{Q}}L_{T}(\Lambda)=\mathbb{E}_{\mathbb{Q}}\Psi(\Lambda)=1$. ∎
By Girsanov’s theorem, we deduce that
Proposition 2.8.
On $\mathrm{d}\mathbb{P}:=L_{T}(\Lambda)\mathrm{d}\mathbb{Q}$, it holds that $(N_{t})$ is a point process of stochastic intensity $f_{\Lambda}(Y_{t-})$. For any measurable test function ϕ:+:italic-$\phi:\mathbb{R}\rightarrow\mathbb{R}_{+}$, consider:
$\langle\phi,M^{\gamma}_{t}\rangle:=\mathbb{E}_{\mathbb{P}}[\phi(\Lambda)~{}|~{}{\mathcal{F}}^{N}_{t}].$
It holds that
$\langle\phi,M^{\gamma}_{t}\rangle$ $=\frac{\mathbb{E}_{\mathbb{Q}}[\phi(\Lambda)L_{t}(\Lambda)~{}|~{}{\mathcal{F}}^{N}_{t}]}{\mathbb{E}_{\mathbb{Q}}[L_{t}(\Lambda)~{}|~{}{\mathcal{F}}^{N}_{t}]}$
$=\frac{\int\varphi(\lambda)L_{t}(\lambda)\mu(\mathrm{d}\lambda)}{\int L_{t}(\lambda)\mu(\mathrm{d}\lambda)}=\frac{\langle\phi L_{t},\mu\rangle}{\langle L_{t},\mu\rangle}.$
Proof.
This follows from [5, Th. 3] and [5, Lem. 5]. For completeness, we give the arguments for the second part of the result. Take $A\in{\mathcal{F}}^{N}_{t}$. By definition of conditional expectation, we have $\mathbb{E}_{\mathbb{P}}[\mathbbm{1}_{A}\phi(\Lambda)]=\mathbb{E}_{\mathbb{P}}[\mathbbm{1}_{A}\mathbb{E}_{\mathbb{P}}[\phi(\Lambda)~{}|~{}{\mathcal{F}}^{N}_{t}]]$. Therefore,
$\mathbb{E}_{\mathbb{Q}}[\mathbbm{1}_{A}\phi(\Lambda)L_{T}(\Lambda)]=\mathbb{E}_{\mathbb{Q}}[\mathbbm{1}_{A}L_{T}(\Lambda)\mathbb{E}_{\mathbb{P}}[\phi(\Lambda)~{}|~{}{\mathcal{F}}^{N}_{t}]].$
Taking the conditional expectation with respect to ${\mathcal{F}}^{N}_{t}$, we obtain:
$\mathbb{E}_{\mathbb{Q}}[\mathbbm{1}_{A}\mathbb{E}_{\mathbb{Q}}[\phi(\Lambda)L_{T}(\Lambda)~{}|~{}{\mathcal{F}}^{N}_{t}]]=\mathbb{E}_{\mathbb{Q}}[\mathbbm{1}_{A}\mathbb{E}_{\mathbb{Q}}[L_{T}(\Lambda)~{}|~{}{\mathcal{F}}^{N}_{t}]\mathbb{E}_{\mathbb{P}}[\phi(\Lambda)~{}|~{}{\mathcal{F}}^{N}_{t}]].$
This is holds for all $A\in{\mathcal{F}}^{N}_{t}$, we find that:
$\mathbb{E}_{\mathbb{Q}}[\phi(\Lambda)L_{T}(\Lambda)~{}|~{}{\mathcal{F}}^{N}_{t}]=\mathbb{E}_{\mathbb{Q}}[L_{T}(\Lambda)~{}|~{}{\mathcal{F}}^{N}_{t}]\mathbb{E}_{\mathbb{P}}[\phi(\Lambda)~{}|~{}{\mathcal{F}}^{N}_{t}],$
giving the first equality for $\langle\phi,M^{\gamma}_{t}\rangle$. For the second equality, note that on $\mathbb{Q}$, $(N_{t})$ and $\Lambda$ are independent, therefore:
$\mathbb{E}_{\mathbb{Q}}[\phi(\Lambda)L_{T}(\Lambda)~{}|~{}{\mathcal{F}}^{N}_{t}]=\int_{\mathbb{R}_{+}}\phi(\lambda)L_{t}(\lambda)\mu(\mathrm{d}\lambda).$
This implies the stated formula. ∎
Corollary 2.9.
Therefore, it holds that $a.s.$,
$M^{\gamma}_{t}:={\mathcal{L}}_{\mathbb{P}}(\Lambda~{}|~{}{\mathcal{F}}^{N}_{t})\propto\prod_{\tau_{i}\leq t}f_{\lambda}(Y_{\tau_{k}-})\exp\left(-\int_{0}^{t}f_{\lambda}(Y_{s})\mathrm{d}s\right)\mu(\mathrm{d}\lambda).$
Following the Bayesian terminology, $M^{\gamma}_{t}$ is the posterior distribution. In addition, the variance of the posterior distribution $V(N):=Var(M^{\gamma}_{T})=\mathbb{E}_{\mathbb{P}}[\Lambda^{2}~{}|~{}{\mathcal{F}}^{N}_{T}]-\left(\mathbb{E}_{\mathbb{P}}(\Lambda~{}|~{}{\mathcal{F}}^{N}_{T})\right)^{2}$ is equal to:
$V(N)=\frac{\langle\phi_{2}L_{t},\mu\rangle}{\langle L_{t},\mu\rangle}-\frac{\langle\phi_{1}L_{t},\mu\rangle}{\langle L_{t},\mu\rangle},\quad\text{ where }\quad\phi_{1}(\lambda)=\lambda\text{ and }\phi_{2}(\lambda)=\lambda^{2}.$
Remark 2.10.
Let ϕ:+:italic-$\phi:\mathbb{R}\rightarrow\mathbb{R}_{+}$ be a non-negative measurable function. Then, $t\mapsto\langle\phi,M^{\gamma}_{t}\rangle=\mathbb{E}_{\mathbb{P}}[\phi(\Lambda)~{}|~{}{\mathcal{F}}^{N}_{t}]$ is a martingale. In other words, $M^{\gamma}_{t}$ is a measure-valued martingale. In particular, it holds that $\mathbb{E}_{\mathbb{P}}\mathbb{E}_{\mathbb{P}}[\Lambda^{2}~{}|~{}{\mathcal{F}}^{N}_{t}]=\mathbb{E}\Lambda^{2}<\infty$. Therefore, for any admissible control $\gamma$, we find that
$\mathbb{P}(\mathrm{d}\omega)a.s.,\quad\mathbb{E}_{\mathbb{P}}[\Lambda^{2}~{}|~{}{\mathcal{F}}^{N}_{t}]<\infty.$
In addition, by the law of the total variance, we have:
$Var(\mu)=Var(M^{\gamma}_{0})=\mathbb{E}_{\mathbb{P}}Var(M^{\gamma}_{t})+Var(\mathbb{E}_{\mathbb{P}}[\Lambda~{}|~{}{\mathcal{F}}^{N}_{t}]).$
This shows in particular that, in expectation, the variance of $M^{\gamma}_{t}$ can only decay with time.
Altogether, we have proven that our optimization problem (2.1) is equal to:
$\inf_{\gamma\in{\mathcal{A}}}\mathbb{E}_{\mathbb{Q}}\left[L_{T}(\Lambda)\left(\int_{0}^{T}\gamma^{2}_{u}\mathrm{d}u+\kappa V(N)\right)\right]$
$=\inf_{\gamma\in{\mathcal{A}}}\int\mu(\mathrm{d}\lambda)\mathbb{E}_{\mathbb{Q}}\left[L_{T}(\lambda)\left(\int_{0}^{T}\gamma^{2}_{u}\mathrm{d}u+\kappa V(N)\right)\right].$
We again use the fact that under $\mathbb{Q}$, the standard Poisson process $N$ of rate $1$ and $\Lambda$ are independent.
2.3 The value function
We now slightly generalize the formulation of the problem to define the value function.
Probability spaces.
We consider the canonical space $(\Omega,{\mathcal{F}},({\mathcal{F}}^{N}_{t}),\mathbb{Q})$ associated to a standard Poisson process $(N_{t})$ with rate $1$. We denote by $(\tau_{k})$ the successive jump times of $(N_{t})$.
Controls.
Denote by ${\mathcal{A}}$ the set of all $({\mathcal{F}}^{N}_{t})$-predictable processes $\gamma$ such that almost surely, $\int_{0}^{T}\gamma^{2}_{u}\mathrm{d}u<\infty$.
Value function.
Given $\gamma\in{\mathcal{A}},y\in\mathbb{R}$ and $\mu\in{\mathcal{P}}_{2}(\mathbb{R}_{+})$, we let:
$J^{\gamma}(t,y,\mu):=\mathbb{E}_{\mathbb{Q}}\left[\int_{\mathbb{R}_{+}}\left(\int_{t}^{T}\frac{\gamma_{s}^{2}}{2}\mathrm{d}s+\text{Var}(M^{t,y,\mu,\gamma}_{T})\right)L^{y,\gamma}_{t,T}(\lambda)\mu(\mathrm{d}\lambda)\right],$
where for all $s\in[t,T]$:
$L^{y,\gamma}_{t,s}(\lambda)=\prod_{\tau_{i}\in(t,s]}f_{\lambda}(Y^{t,y,\gamma}_{\tau_{i}-})\exp\left((s-t)-\int_{t}^{s}f_{\lambda}(Y^{t,y,\gamma}_{u})\mathrm{d}u\right),$ (2.2)
$Y^{t,y,\gamma}_{s}=y+\int_{t}^{s}\gamma_{u}\mathrm{d}u-\int_{t}^{s}Y^{t,y,\gamma}_{u-}\mathrm{d}N_{u},$ (2.3)
and the posterior measure $M^{t,y,\mu,\gamma}_{s}$ is defined by:
$\langle\phi,M^{t,y,\mu,\gamma}_{s}\rangle:=\frac{\langle\phi L^{y,\gamma}_{t,s},\mu\rangle}{\langle L^{y,\gamma}_{t,s},\mu\rangle},$ (2.4)
this equality being true for any non-negative measurable function ϕ:++:italic-$\phi:\mathbb{R}_{+}\rightarrow\mathbb{R}_{+}$. Moreover, the variance of $M^{t,y,\mu,\gamma}_{T}$ is:
$\text{Var}(M^{t,y,\mu,\gamma}_{T})=\langle\phi_{2},M^{t,y,\mu,\gamma}_{T}\rangle-(\langle\phi_{1},M^{t,y,\mu,\gamma}_{T}\rangle)^{2},\quad\text{ with }\phi_{1}(\lambda)=\lambda,\phi_{2}(\lambda)=\lambda^{2}.$
Finally, the value function is
$v(t,y,\mu):=\inf_{\gamma\in{\mathcal{A}}}J^{\gamma}(t,x,\mu).$
We have proven in the previous sections that:
Theorem 2.11.
The value define by the optimization problem (2.1) is equal to $v(0,y,\mu)$.
2.4 Reduction to finite dimension
In what follows, we assume that:
Assumption 2.12.
There is a globally Lipschitz and bounded function $g:\mathbb{R}\rightarrow\mathbb{R}_{+}$ such that:
$\forall y\in\mathbb{R},\lambda\in\mathbb{R}_{+},\quad f_{\lambda}(y)=\lambda g(y).$
Surprisingly, with this assumption, the optimization problem over the space of probability measures can be reduced to a finite dimension problem. To proceed, given $\mu\in{\mathcal{P}}_{2}(\mathbb{R}_{+})$, we set for all $z\geq 0$ and $n\in\mathbb{N}$:
$\tilde{v}(t,x,z,n):=v(t,x,m_{\mu}(n,z)),\quad\text{ where }\quad m_{\mu}(n,z)(\mathrm{d}\lambda):=\frac{\lambda^{n}e^{-\lambda z}\mu(\mathrm{d}\lambda)}{\int_{\mathbb{R}_{+}}\theta^{n}e^{-\theta z}\mu(\mathrm{d}\theta)}.$
Note that $m_{\mu}(0,0)=\mu$. We also set
$\Phi_{\mu}(n,z):=\int_{\mathbb{R}_{+}}\lambda^{n}e^{-\lambda z}\mu(\mathrm{d}\lambda),$
and note that
$\forall k\in\mathbb{N},\quad\int_{\mathbb{R}_{+}}\lambda^{k}m_{\mu}(n,z)(\mathrm{d}\lambda)=\frac{\Phi_{\mu}(n+k,z)}{\Phi_{\mu}(n,z)}.$
Finally, we let:
$\Psi_{\mu}(n,z):=\text{Var}(m_{\mu}(n,z))=\frac{\Phi_{\mu}(n+2,z)}{\Phi_{\mu}(n,z)}-\left(\frac{\Phi_{\mu}(n+1,z)}{\Phi_{\mu}(n,z)}\right)^{2}.$ (2.5)
Recall the definition of $M^{t,y,\mu,\gamma}_{s}$ and $Y^{t,y,\gamma}_{s}$ given by equations (2.4) and  (2.3). We define:
$Z^{t,y,z,\gamma}_{s}:=z+\int_{t}^{s}g(Y^{t,y,\gamma}_{u})\mathrm{d}u.$ (2.6)
Lemma 2.13.
Let $\gamma\in{\mathcal{A}}$. It holds that, almost surely:
$M^{t,y,m_{\mu}(n,z),\gamma}_{s}=m_{\mu}(n+N_{s}-N_{t},Z^{t,y,z,\gamma}_{s}),$
Proof.
This follows from the fact that for every measurable non-negative function $\phi$, we have
$\langle\phi,m_{\mu}(n+N_{s}-N_{t},Z^{t,y,z,\gamma}_{s})\rangle=\frac{\langle\phi L^{y,\gamma}_{t,s},m_{\mu}(n,z)\rangle}{\langle L^{y,\gamma}_{t,s},m_{\mu}(n,z)\rangle}.$ (2.7)
As a corollary, we obtain:
Proposition 2.14.
It holds that
$\tilde{v}(t,y,z,n)=\inf_{\gamma\in{\mathcal{A}}}\mathbb{E}_{\mathbb{Q}}\left[\int_{\mathbb{R}_{+}}\left(\int_{t}^{T}\frac{\gamma_{s}^{2}}{2}\mathrm{d}s+\Psi_{\mu}(n+N_{T}-N_{t},Z^{t,y,z,\gamma}_{T})\right)L^{t,y,\gamma}_{T}(\lambda)m_{\mu}(n,z)(\mathrm{d}\lambda)\right].$
The key point is that this optimization problem is now a classical stochastic optimization in finite dimension. For now on, we will concentrate our study on this finite-dimensional problem. Therefore, we write to simplify the notations $v(t,y,z,n)=\tilde{v}(t,y,z,n)$.
Some properties of the posterior distribution
Finally, we close this section with some properties satisfied by $m_{\mu}(n,z)$.
Lemma 2.15.
Let $\Xi_{\mu}(n,z):=\Phi_{\mu}(n+1,z)/\Phi_{\mu}(n,z)$ be the expected value of the posterior distribution. We have for all $n\in\mathbb{N}$ and $z\geq 0$:
  1. 1.
    $\partial_{z}\Phi_{\mu}(n,z)=-\Phi_{\mu}(n+1,z)$.
  2. 2.
    $\partial_{z}\Xi_{\mu}(n,z)=-\Psi_{\mu}(n,z)$.
  3. 3.
    $\Xi_{\mu}(n+1,z)-\Xi_{\mu}(n,z)=\frac{\Psi_{\mu}(n,z)}{\Xi_{\mu}(n,z)}$.
  4. 4.
    $\partial_{z}\Psi(n,z)=\Xi_{\mu}(n,z)\left[\Psi_{\mu}(n,z)-\Psi_{\mu}(n+1,z)-\left(\frac{\Psi_{\mu}(n,z)}{\Xi_{\mu}(n,z)}\right)^{2}\right]$.
Remark 2.16.
The points 2 and 3 above show that the mean of posterior distribution decays between the jumps, while it always increase just after a jump. However, a similar behavior does not hold for the variance of the posterior distribution. Consider for instance the case where $\mu(\mathrm{d}\lambda)=\frac{1}{2}\delta_{0}+\frac{1}{2}\delta_{\lambda_{\max}}$. After one jump, the posterior distribution is equal to $\delta_{\lambda_{\max}}$, with variance zero: $n\geq 1\implies\Psi_{\mu}(n,z)=0$. In that case, the variance decays to zero just after the first jump.
Lemma 2.17.
Assume that $\text{Supp}(\mu)\subset[0,\lambda_{\max}]$. Then $\Psi_{\mu}(n,z)\leq\frac{\lambda^{2}_{\max}}{4}$ and
$\lvert\partial_{z}\Psi_{\mu}(n,z)\rvert\leq\frac{1}{2}\lambda^{3}_{\max}.$
3 Dynamic programming principle and regularity
3.1 Space regularity
We now study the regularity of the value function with respect to $y$ and $z$. Recall that
$v(t,y,z,n)=\inf_{\gamma\in{\mathcal{A}}}J_{1}^{\gamma}(t,y,z,n)+J^{\gamma}_{2}(t,y,z,n),$
with
$J_{1}^{\gamma}(t,y,z,n)$ $:=\mathbb{E}_{\mathbb{Q}}\left[L^{\gamma,y}_{t,T}(\Lambda)\int_{t}^{T}\gamma^{2}_{s}\mathrm{d}s\right],$
$J_{2}^{\gamma}(t,y,z,n)$ $:=\mathbb{E}_{\mathbb{Q}}\left[L^{\gamma,y}_{t,T}(\Lambda)\Psi_{\mu}(n+N_{T}-N_{t},z+\int_{t}^{T}g(Y^{t,y,\gamma}_{s})\mathrm{d}s)\right],$
where under $\mathbb{Q}$, $\Lambda$ is a random variable with law $\mathbb{P}(\Lambda\in\mathrm{d}\lambda)=\lambda^{n}e^{-\lambda z}\mu(\mathrm{d}\lambda)/\Phi_{\mu}(n,z)$ and $(N_{t})$ is a standard Poisson process with rate $1$ independent of $\Lambda$.
Assumption 3.1.
We assume that
  1. 1.
    The prior distribution is compactly supported: $\exists\lambda_{\max}>0$:
    $\text{Supp}(\mu)\subset[0,\lambda_{\max}].$
  2. 2.
    The function $g\in C^{1}(\mathbb{R};\mathbb{R}_{+})$ with $\lVert g\rVert_{\infty}+\lVert\frac{g^{\prime}}{g}\rVert_{\infty}<\infty$.
  3. 3.
    The function $b\in C^{1}(\mathbb{R})$ with $\lVert b^{\prime}\rVert_{\infty}<\infty$.
The main result of this section is the following.
Proposition 3.2.
Under Assumption 3.1, it holds that there exists a constant $C_{T}$ such that for all $n\geq 0$ and for all $x,x^{\prime}\in\mathbb{R}$ and all $z,z^{\prime}\geq 0$:
$\lvert v(t,y,z,n)-v(t,y^{\prime},z^{\prime},n)\rvert\leq C_{T}(|y-y^{\prime}|+|z-z^{\prime}|).$
We now give the proof of this result. To simplify the notations, we assume that $t=0$ and write:
$L^{y,\gamma}_{T}(\lambda)=L^{y,\gamma}_{t,T}(\lambda).$
3.1.1 Regularity with respect to $y$
Lemma 3.3.
There exists a constant $C_{T}$ such that for all $\gamma\in{\mathcal{A}}$, it holds that:
$\forall y\in\mathbb{R},\quad\lvert\partial_{y}L^{y,\gamma}_{T}(\lambda)\rvert\leq C_{T}L^{y,\gamma}_{T}(\lambda).$
Proof.
We prove the result with $C_{T}:=\left(\lVert\frac{g^{\prime}}{g}\rVert_{\infty}+\lambda_{\max}T\lVert g^{\prime}\rVert_{\infty}\right)e^{T\lVert b^{\prime}\rVert_{\infty}}$. Observe that just after a jump, $Y$ is reset to zero and so the initial condition $y$ is forgotten. We write $\varphi_{s}(y)$ for the solution of the ODE $\frac{d}{ds}\varphi_{s}(y)=b(\varphi_{s}(y))+\gamma_{u}$ with initial condition $\varphi_{0}(y)=y$. Assume first that $N_{T}=0$. Then
$\partial_{y}L^{y,\gamma}_{T}(\lambda)=-\lambda L^{y,\gamma}_{T}(\lambda)\int_{0}^{T}g^{\prime}(\varphi_{s}(y))\partial_{y}\varphi_{s}(y)\mathrm{d}s.$
As $\partial_{y}\varphi_{s}(y)=\exp\left(\int_{0}^{s}b^{\prime}(\varphi_{u}(y))\mathrm{d}u\right)\leq\exp(T\lVert b^{\prime}\rVert_{\infty})$, the result holds.
Assume now that $N_{T}>0$. Then, provided that $\tau_{1}$ is the time of the first jump of $N$, we have:
$\partial_{y}L^{y,\gamma}_{T}(\lambda)=\left[\frac{g^{\prime}(\varphi_{\tau_{1}-}(y))\partial_{y}\varphi_{\tau_{1}-}(y)}{g(\varphi_{\tau_{1}-}(y))}-\lambda\int_{0}^{\tau_{1}}g^{\prime}(\varphi_{s}(y))\partial_{y}\varphi_{s}(y)\mathrm{d}s\right]L^{y,\gamma}_{T}(\lambda).$
We deduce the result using our assumptions on $g$ and $b$. ∎
Lemma 3.4.
There exists a constant $C_{T}$ such that for all $\gamma\in{\mathcal{A}}$,
$\forall y,y^{\prime}\in\mathbb{R},\quad L^{y^{\prime},\gamma}_{T}(\lambda)\leq e^{C_{T}|y-y^{\prime}|}L^{y,\gamma}_{T}(\lambda).$
Proof.
Assume first that $y^{\prime}>y$. Then the result follows by Grönwall’s lemma. If now $y^{\prime}<y$, we set $\phi(s)=L^{y-s,\gamma}_{T}(\lambda)$. We have $\phi^{\prime}(s)=-\partial_{y}L^{y-s,\gamma}_{T}(\lambda)\leq C_{T}\phi(s)$. Therefore, by Grönwall’s lemma, $\phi(s)\leq e^{C_{T}s}\phi(0).$ Choosing $s=y-y^{\prime}$ ends the proof. ∎
Lemma 3.5.
There is another constant $C_{T}$ such that for any $\gamma\in{\mathcal{A}}$, for all $y,y^{\prime}$ with $|y-y^{\prime}|\leq 1$:
$\lvert L^{y,\gamma}_{T}(\lambda)-L^{y^{\prime},\gamma}_{T}(\lambda)\rvert\leq C_{T}|y-y^{\prime}|(L^{y,\gamma}_{T}(\lambda)+L^{y^{\prime},\gamma}_{T}(\lambda)).$
Proof.
Without loss of generality, assume that $y+1>y^{\prime}>y$. Then we have $L^{y^{\prime},\gamma}_{T}(\lambda)-L^{y,\gamma}_{T}(\lambda)\leq(e^{C_{T}(y^{\prime}-y)}-1)L^{y,\gamma}_{T}(\lambda)\leq C_{T}e^{C_{T}}|y-y^{\prime}|L^{y,\gamma}_{T}(\lambda)$. Similarly, $L^{y,\gamma}_{T}(\lambda)-L^{y^{\prime},\gamma}_{T}(\lambda)\leq C_{T}e^{C_{T}}|y-y^{\prime}|L^{y^{\prime},\gamma}_{T}(\lambda)$. Altogether, the result holds. ∎
Corollary 3.6.
As $L^{y^{\prime},\gamma}_{T}(\lambda)\leq e^{C_{T}}L^{y,\gamma}_{T}(\lambda)$, we deduce that there exists another constant $C_{T}$ such that for all $y,y^{\prime}$ with $|y-y^{\prime}|\leq 1$,
$\lvert L^{y,\gamma}_{T}(\lambda)-L^{y^{\prime},\gamma}_{T}(\lambda)\rvert\leq C_{T}|y-y^{\prime}|L^{y,\gamma}_{T}(\lambda).$
Lemma 3.7.
There exists a constant $C_{T}$ such that for any $\gamma\in{\mathcal{A}}$, it holds that for all $y,y^{\prime}$ with $|y-y^{\prime}|\leq 1$:
$\lvert J^{\gamma}_{1}(t,y^{\prime},z,n)-J_{1}^{\gamma}(t,y,z,n)\rvert$ $\leq C_{T}|y-y^{\prime}|J^{\gamma}_{1}(t,y,z,n),$
$\lvert J_{2}^{\gamma}(t,y^{\prime},z,n)-J_{2}^{\gamma}(t,y,z,n)\rvert$ $\leq C_{T}|y-y^{\prime}|(1+J^{\gamma}_{2}(t,y,z,n)).$
Proof.
We have
$\lvert J^{\gamma}_{1}(t,y^{\prime},z,n)-J^{\gamma}_{1}(t,y,z,n)\rvert$ $\leq\mathbb{E}_{\mathbb{Q}}\left[\lvert L^{y^{\prime},\gamma}_{t,T}(\Lambda)-L^{y,\gamma}_{t,T}(\Lambda)\rvert\int_{t}^{T}\frac{\gamma_{s}^{2}}{2}\mathrm{d}s\right]$
$\leq C_{T}|y-y^{\prime}|\mathbb{E}_{\mathbb{Q}}\left[L^{y,\gamma}_{t,T}(\Lambda)\int_{t}^{T}\frac{\gamma_{s}^{2}}{2}\mathrm{d}s\right]=C_{T}|y-y^{\prime}|J^{\gamma}_{1}(t,y,z,n).$
Similarly,
$\lvert J^{\gamma}_{2}(t,y,z,n)-J^{\gamma}_{2}(t,y^{\prime},z,n)\rvert\leq$
$\quad\mathbb{E}_{\mathbb{Q}}\lvert L^{\gamma,y}_{T}(\Lambda)-L^{\gamma,y^{\prime}}_{T}(\Lambda)\rvert\Psi_{\mu}(n+N_{T}-N_{t},Z^{t,y,z,\gamma}_{T})$
$\quad+\mathbb{E}_{\mathbb{Q}}\left[L^{\gamma,y^{\prime}}_{T}(\Lambda)\left(\Psi_{\mu}(n+N_{T}-N_{t},Z^{t,y,z,\gamma}_{T})-\Psi_{\mu}(n+N_{T}-N_{t},Z^{t,y^{\prime},z,\gamma}_{T})\right)\right]=:A+B.$
We have $A\leq C_{T}|y-y^{\prime}|\mathbb{E}_{\mathbb{Q}}L^{\gamma,y}_{t,T}(\Lambda)\Psi_{\mu}(n+N_{T}-N_{t},Z^{t,y,z,\gamma}_{T})=C_{T}|y-y^{\prime}|J^{\gamma}_{2}(t,y,z,n)$ and:
$B$ $\leq\mathbb{E}_{\mathbb{Q}}L^{y^{\prime},\gamma}_{t,T}(\Lambda)\lVert\partial_{z}\Psi_{\mu}(n+N_{T}-N_{t},\cdot)\rVert_{\infty}\int_{t}^{T}\lVert g^{\prime}\rVert_{\infty}e^{\lVert b^{\prime}\rVert_{\infty}T}|y-y^{\prime}|\mathrm{d}s$
$\leq C_{T}\mathbb{E}_{\mathbb{Q}}L^{y,\gamma}_{t,T}(\Lambda)|y-y^{\prime}|=C_{T}|y-y^{\prime}|.$
We used that $\sup_{n\in\mathbb{N}}\lVert\partial_{z}\Psi_{\mu}(n,\cdot)\rVert<\infty$. We deduce the result. ∎
Altogether, we deduce that
$\lvert J^{\gamma}(t,y,z,n)-J^{\gamma}(t,y^{\prime},z,n)\rvert\leq C_{T}(1+J^{\gamma}(t,y,z,n))|y-y^{\prime}|.$
Let $y,y^{\prime}\in\mathbb{R}$ be fixed. Consider $\epsilon>0$ and $\gamma$ an $\epsilon$-optimal control in the sense that
$v(t,y,z,n)\geq J^{\gamma}(t,y,z,n)-\epsilon.$
We have
$v(t,y^{\prime},z,n)-v(t,y,z,n)$ $\leq J^{\gamma}(t,y^{\prime},z,n)-J^{\gamma}(t,y,z,n)+\epsilon$
$\leq C_{T}|y-y^{\prime}|(1+J^{\gamma}(t,t,z,n))+\epsilon$
$\leq C_{T}|y-y^{\prime}|(1+\epsilon+v(t,y,z,n))+\epsilon.$
Sending $\epsilon\downarrow 0$, we deduce that
$v(t,y^{\prime},z,n)-v(t,y,z,n)\leq C_{T}|y^{\prime}-y|(1+v(t,y,z,n)).$
In addition, note that
$v(t,y,z,n)\leq J^{0}(t,y,z,n)\leq\frac{\lambda^{2}_{\max}}{4}<\infty.$
Exchanging the role of $y$ and $y^{\prime}$ we deduce that
$\lvert v(t,y,z,n)-v(t,y^{\prime},z,n)\rvert\leq C_{T}|y-y^{\prime}|.$
3.1.2 Regularity with respect to $z$
The regularity with respect to $z$ is proven by similar arguments. Recall that
$\Phi_{\mu}(n,z)=\int_{\mathbb{R}_{+}}\lambda^{n}e^{-\lambda z}\mu(\mathrm{d}\lambda),\quad m_{\mu}(n,z)=\frac{\lambda^{n}e^{-\lambda z}}{\Phi_{\mu}(n,z)}\mu(\mathrm{d}\lambda).$
We have
Lemma 3.8.
There exists constants $C,\eta>0$ such that for all $n\in\mathbb{N}$:
  1. 1.
    $\forall z,z^{\prime}\geq 0$, $\lvert\Phi_{\mu}(n,z)-\Phi_{\mu}(n,z^{\prime})\rvert\leq C|z-z^{\prime}|\Phi_{\mu}(n,z)$.
  2. 2.
    $\forall z,z^{\prime}\geq 0$, $\lvert m_{\mu}(n,z)-m_{\mu}(n,z^{\prime})\rvert\leq C|z-z^{\prime}|[m_{\mu}(n,z)+m_{\mu}(n,z^{\prime})]$.
  3. 3.
    $\forall z,z^{\prime}\geq 0$ with $|z-z^{\prime}|<\eta$, $\lvert m_{\mu}(n,z)-m_{\mu}(n,z^{\prime})\rvert\leq C|z-z^{\prime}|m_{\mu}(n,z)$.
Proof.
We start with the first point. We have
$\lvert\Phi_{\mu}(n,z)-\Phi_{\mu}(n,z^{\prime})\rvert\leq\int_{\mathbb{R}_{+}}\lambda^{n}e^{-\lambda z}\lvert 1-e^{-\lambda(z^{\prime}-z)}\rvert\mu(\mathrm{d}\lambda).$
As $Supp(\mu)\subset[0,\lambda_{\max}]$, the result holds. For the second point, we have
$\lvert m_{\mu}(n,z)-m_{\mu}(n,z^{\prime})\rvert\leq\left[\frac{\lvert\Phi_{\mu}(n,z^{\prime})-\Phi_{\mu}(n,z)\rvert e^{-\lambda z}}{\Phi_{\mu}(n,z)\Phi_{\mu}(n,z^{\prime})}+\frac{e^{-\lambda z^{\prime}}(e^{-\lambda(z-z^{\prime})}-1)}{\Phi(n,z^{\prime})}\right]\lambda^{n}\mu(\mathrm{d}\lambda).$
Using the first point, we deduce that the second inequality holds. Finally, we verify that provided $|z-z^{\prime}|\leq\eta$, $\eta$ small enough, it holds that
$m_{\mu}(n,z^{\prime})\leq Cm_{\mu}(n,z).$
Indeed, we have
$m_{\mu}(n,z^{\prime})=\frac{\lambda^{n}e^{-\lambda z}e^{-\lambda(z^{\prime}-z)}\mu(\mathrm{d}\lambda)}{\Phi_{\mu}(n,z^{\prime})-\Phi_{\mu}(n,z)+\Phi_{\mu}(n,z)}.$
As $\lvert\Phi_{\mu}(n,z^{\prime})-\Phi_{\mu}(n,z)\rvert\leq\frac{1}{2}\Phi_{\mu}(n,z)$ provided that $\lvert z-z^{\prime}\rvert$ is small enough, we deduce the last statement. ∎
As on $\mathbb{Q}$, the Poisson process $N$ is independent of $\Lambda$, we have
$J^{\gamma}_{1}(t,y,z,n)=\int_{\mathbb{R}_{+}}\left[\mathbb{E}_{\mathbb{Q}}L^{y,\gamma}_{t,T}(\lambda)\int_{t}^{T}\frac{\gamma^{2}_{s}}{2}\mathrm{d}s\right]m_{\mu}(n,z)(\mathrm{d}\lambda).$
We deduce from the last point of the previous Lemma that for all $z,z^{\prime}\geq 0$, $\lvert z-z^{\prime}\rvert\leq\eta$,
$\lvert J^{\gamma}_{1}(t,y,z,n)-J^{\gamma}_{1}(t,y,z^{\prime},n)\rvert\leq C_{T}\lvert z-z^{\prime}\rvert J^{\gamma}_{1}(t,y,z,n).$
Similarly, we have
$\lvert J^{\gamma}_{2}(t,y,z,n)-J^{\gamma}_{2}(t,y,z^{\prime},n)\rvert\leq C_{T}|z-z^{\prime}|(1+J^{\gamma}_{2}(t,y,z,n)).$
We deduce as previously that
$\lvert v(t,y,z,n)-v(t,y,z^{\prime},n)\rvert\leq C_{T}\lvert z-z^{\prime}\rvert.$
This ends the proof of Proposition 3.2.
3.2 A first dynamic programming principle
Recall that the value function is defined by
$v(t,y,z,n)$ $:=\inf_{\gamma\in{\mathcal{A}}}\mathbb{E}_{\mathbb{Q}}\langle L^{y,\gamma}_{t,T},m_{\mu}(n,z)\rangle\left[\int_{t}^{T}\frac{\gamma^{2}_{u}}{2}\mathrm{d}u+\kappa\Psi_{\mu}(n+N_{T}-N_{t},Z_{T}^{t,y,z,\gamma})\right]$
$=\inf_{\gamma\in{\mathcal{A}}}J^{\gamma}(t,y,z,n).$
[…]
Proposition 3.9.
It holds that for all $s\in[t,T]$,
$v(t,y,z,n)=\inf_{\gamma\in{\mathcal{A}}}$ $\mathbb{E}_{\mathbb{Q}}\langle L^{y,\gamma}_{t,s},m_{\mu}(n,z)\rangle\left[\int_{t}^{s}\frac{\gamma^{2}_{u}}{2}\mathrm{d}u+v(s,Y^{t,y,\gamma}_{s},Z^{t,y,z,\gamma}_{s},n+N_{s}-N_{t})\right].$
We now detail the proof of this result. Recall that for $\gamma\in{\mathcal{A}}$,
$J^{\gamma}_{1}(t,y,z,n)$ $=\mathbb{E}_{\mathbb{Q}}\left[\langle L^{\gamma,y}_{t,T},m_{\mu}(n,z)\rangle\int_{t}^{T}\frac{\gamma^{2}_{u}}{2}\mathrm{d}u\right],$
$J^{\gamma}_{2}(t,y,z,n)$ $=\mathbb{E}_{\mathbb{Q}}\left[\langle L^{\gamma,y}_{t,T},m_{\mu}(n,z)\rangle\Psi_{\mu}(n+N_{T}-N_{t},Z^{t,y,z,\gamma}_{T})\right].$
We start with the following key lemma.
Lemma 3.10.
Let $S$ be a $({\mathcal{F}}^{N}_{t})$-stopping time such that $a.s.$, $S\in[t,T]$. It holds that $a.s$,
$\mathbb{E}_{\mathbb{Q}}\left[\langle L^{y,\gamma}_{t,T},m_{\mu}(n,z)\rangle\int_{S}^{T}\frac{\gamma^{2}_{u}}{2}\mathrm{d}u\mid{\mathcal{F}}^{N}_{S}\right]$
$\quad=\langle L^{y,\gamma}_{t,S},m_{\mu}(n,z)\rangle J^{\gamma}_{1}(S,Y^{t,y,\gamma}_{S},Z^{t,y,z,\gamma}_{S},n+N_{S}-N_{t}),$
and
$\mathbb{E}_{\mathbb{Q}}[\langle L^{y,\gamma}_{t,T},m_{\mu}(n,z)\rangle\Psi_{\mu}(n+N_{T}-N_{t},Z^{t,y,z,\gamma}_{T})~{}|~{}{\mathcal{F}}^{N}_{S}]$
$\quad=\langle L^{y,\gamma}_{t,S},m_{\mu}(n,z)\rangle J^{\gamma}_{2}(S,Y^{t,y,\gamma}_{S},Z^{t,y,z,\gamma}_{S},n+N_{S}-N_{t}).$
Proof.
First, we note that for all $\lambda\geq 0$:
$a.s,\quad L^{y,\gamma}_{t,T}(\lambda)=L^{y,\gamma}_{t,S}(\lambda)L^{\gamma,Y^{t,y,\gamma}_{S}}_{S,T}(\lambda).$
We deduce from (2.7) that
$\langle L^{y,\gamma}_{t,S},m_{\mu}(n,z)\rangle\langle L^{Y^{t,y,\gamma}_{S},\gamma}_{S,T},m_{\mu}(n+N_{S}-N_{t},Z^{t,y,z,\gamma}_{S}\rangle=\langle L^{y,\gamma}_{t,T},m_{\mu}(n,z)\rangle.$
We now give the proof of the second equality. Take $A\in{\mathcal{F}}^{N}_{S}$. To simplify the notations, write:
$y_{S}:=Y^{t,y,\gamma}_{S},\quad z_{S}:=Z^{t,y,z,\gamma}_{S},\quad n_{S}:=n+N_{S}-N_{t}.$
We also set:
$\Delta Z_{S,T}:=Z_{T}^{S,y_{S},z_{S},\gamma}-z_{S}=\int_{S}^{T}g(Y^{t,y,\gamma}_{u})\mathrm{d}u.$
We have
$\mathbb{E}_{\mathbb{Q}}\langle L^{y,\gamma}_{t,T},m_{\mu}(n,z)\rangle\mathbbm{1}_{A}\Psi_{\mu}(n+N_{T}-N_{t},Z^{t,y,z,\gamma}_{T})$
$\quad=\mathbb{E}_{\mathbb{Q}}\langle L^{y,\gamma}_{t,S},m_{\mu}(n,z)\rangle\langle L^{y_{S},\gamma}_{S,T},m_{\mu}(n_{S},z_{S})\rangle\mathbbm{1}_{A}\Psi_{\mu}(n_{S}+N_{T}-N_{S},z_{S}+\Delta Z_{S,T})$
$\quad=\mathbb{E}_{\mathbb{Q}}\langle L^{y,\gamma}_{t,S},m_{\mu}(n,z)\rangle\mathbbm{1}_{A}$
$\qquad~{}~{}~{}~{}~{}\mathbb{E}_{\mathbb{Q}}\left[\langle L^{y_{S},\gamma}_{S,T},m_{\mu}(n_{S},z_{S})\rangle\Psi_{\mu}(n_{S}+N_{T}-N_{S},z_{S}+\Delta Z_{S,T})~{}|~{}{\mathcal{F}}^{N}_{S}\right]$
$\quad=\mathbb{E}_{\mathbb{Q}}\langle L^{y,\gamma}_{t,S},m_{\mu}(n,z)\rangle\mathbbm{1}_{A}J^{\gamma}_{2}(S,y_{S},z_{S},n_{S}).$
The obtain the last identity, we used the strong Markov property satisfied by $(Y^{t,y,\gamma}_{s},n+N_{s}-N_{t},Z^{t,y,z,\gamma}_{s})$. As the equality is valid for any $A\in{\mathcal{F}}^{N}_{S}$, the result is proved. ∎
We now detail the proof of Proposition 3.9, which follow from standard arguments. Recall that $J^{\gamma}(t,y,z,n)=J^{\gamma}_{1}(t,y,z,n)+\kappa J^{\gamma}_{2}(t,y,z,n)$. Using Lemma 3.10 with a deterministic $S=s\in[t,T]$, we have
$v(t,y,z,n)=\inf_{\gamma\in{\mathcal{A}}}\mathbb{E}_{\mathbb{Q}}\left[\langle L^{y,\gamma}_{t,s},m_{\mu}(n,z)\rangle\left\{\int_{t}^{s}\frac{\gamma^{2}_{u}}{2}\mathrm{d}u+J^{\gamma}(s,Y^{t,y,\gamma}_{s},Z^{t,y,z,\gamma}_{s},n+N_{s}-N_{t})\right\}\right].$
We deduce that
$v(t,y,z,n)\geq\inf_{\gamma\in{\mathcal{A}}}\mathbb{E}_{\mathbb{Q}}\left[\langle L^{x,\gamma}_{t,s},m_{\mu}(n,z)\rangle\left\{\int_{t}^{s}\frac{\gamma^{2}_{u}}{2}\mathrm{d}u+v(s,Y^{t,y,\gamma}_{s},Z^{t,y,z,\gamma}_{s},n+N_{s}-N_{t})\right\}\right].$
For the other inequality, fix $\epsilon>0$ and $s\in(0,T)$. Using the regularity of $J^{\gamma}$ and $v$ proven in Section 3.1, there exists $\alpha>0$ such that for all $n\in\mathbb{N}$, there is a partition $(B_{i}^{n})_{i\in\mathbb{N}}$ of $\mathbb{R}\times\mathbb{R}_{+}$ with centers $(y_{i}^{n},z_{i}^{n})\in B_{i}^{n}$ such that: for all $(n,i)\in\mathbb{N}^{2}$, for all $(y,z)\in B_{i}^{n}$,
$|y-y_{i}^{n}|+|z-z_{i}^{n}|<\alpha\quad\text{and}\quad|v(s,y,z,n)-v(s,y_{i}^{n},z_{i}^{n},n)|<\epsilon.$
For each $(y_{i}^{n},z_{i}^{n},n)$, there exists an $\epsilon$-optimal control $\alpha(y_{i}^{n},z_{i}^{n},n)\in{\mathcal{A}}$:
$J^{\alpha(y_{i}^{n},z_{i}^{n},n)}(s,y^{n}_{i},z^{n}_{i},n)\leq v(s,y^{n}_{i},z^{n}_{i},n)+\epsilon.$
Finally, we set for all $\gamma\in{\mathcal{A}}$:
$\gamma^{*}_{u}=\begin{cases}\gamma_{u}\quad\text{ if }\quad u<s\\ \sum_{i\geq 0}\alpha_{u}^{i,n_{s}}\mathbbm{1}_{B^{n_{s}}_{i}}(Y^{t,y,\gamma}_{s},Z^{t,y,z,\gamma}_{s})\quad\text{ if }\quad u\in[s,T].\end{cases}$
where $\alpha^{i,n}:=\alpha(y_{i}^{n},z_{i}^{n},n)$ for $i,n\geq 0$ and $n_{s}:=n+N_{s}-N_{t}$. To simplify the notations, we also write $y_{s}=Y^{t,y,\gamma}_{s}$ and $z_{s}=Z^{t,y,z,\gamma}_{s}$. We have:
$J^{\gamma^{*}}(s,y_{s},z_{s},n_{s})$ $=\sum_{i}\left[J^{\gamma^{*}}(s,y_{s},z_{s},n_{s})-J^{\gamma^{*}}(s,y^{n_{s}}_{i},z^{n_{s}}_{i},n_{s})\right]\mathbbm{1}_{B^{n_{s}}_{i}}(y_{s},z_{s})$
$\quad+\sum_{i}\left[J^{\gamma^{*}}(s,y^{n_{s}}_{i},z^{n_{s}}_{i},n_{s})-v(s,y^{n_{s}}_{i},z^{n_{s}}_{i},n_{s})\right]\mathbbm{1}_{B^{n_{s}}_{i}}(y_{s},z_{s})$
$\quad+\sum_{i}\left[v(s,y^{n_{s}}_{i},z^{n_{s}}_{i},n_{s})-v(s,y_{s},z_{s},n_{s})\right]\mathbbm{1}_{B^{n_{s}}_{i}}(y_{s},z_{s})$
$\quad+v(s,y_{s},z_{s},n_{s})$
$\leq 3\epsilon+v(s,y_{s},z_{s},n_{s}).$
Therefore, we find that
$v(t,y,z,n)\leq 3\epsilon+\mathbb{E}_{\mathbb{Q}}\left[\langle L^{y,\gamma}_{t,s},m_{\mu}(n,z)\rangle\left\{\int_{t}^{s}\frac{\gamma^{2}_{u}}{2}\mathrm{d}u+v(s,y_{s},z_{s},n_{s})\right\}\right].$
We send $\epsilon$ to zero. Using moreover that the control $\gamma$ is arbitrary between $t$ and $s$, we deduce that
$v(t,y,z,n)\leq\inf_{\gamma\in{\mathcal{A}}}\mathbb{E}_{\mathbb{Q}}\left[\langle L^{y,\gamma}_{t,s},m_{\mu}(n,z)\rangle\left\{\int_{t}^{s}\frac{\gamma^{2}_{u}}{2}\mathrm{d}u+v(s,y_{s},z_{s},n_{s})\right\}\right].$
This completes the proof of Proposition 3.9. ∎
3.3 Time regularity
We now prove:
Proposition 3.11.
Under Assumption 3.1, the value function is continuous.
In view of Proposition 3.2, all we have to prove is that the value function is continuous with respect to time. As it is assumed that the prior distribution is compactly supported in $[0,\lambda_{\max}]$, we recall that
$v(t,y,z,n)\leq\frac{\lambda^{2}_{\max}}{4}.$
We start with a lemma. Let $\gamma\in{\mathcal{A}}$. We denote by $\varphi^{\gamma}_{t,s}(y)$ the solution of the ODE
$\frac{\mathrm{d}}{\mathrm{d}s}\varphi^{\gamma}_{t,s}(y)=b(\varphi^{\gamma}_{t,s}(y))+\gamma_{s}\quad\text{ with }\quad\varphi^{\gamma}_{t,t}(y)=y.$
Lemma 3.12.
There exists a constant $C_{T}>0$ such that for all $t\leq s\leq T$, it holds that
$\forall y\in\mathbb{R},\quad\lvert\varphi^{\gamma}_{t,s}(y)-y\rvert\leq C_{T}\int_{t}^{s}\lvert\gamma_{u}\rvert\mathrm{d}u.$
Proof.
This follows from the Grönwall’s lemma applied to $s\mapsto\varphi^{\gamma}_{t,s}(y)-y$, using that $b$ is Lipschitz. ∎
Assume without loss of generalities that $s>t$ and that $s-t\leq 1$. To simplify the notations, we write $Y_{s}=Y^{t,y,\gamma}_{s},n_{s}=n+N_{s}-N_{t}$, $z_{s}=z+\int_{t}^{s}g(Y^{t,y,\gamma}_{u}\mathrm{d}u)$. Using Proposition 3.9,
$v(t,y,z,n)-v(s,y,z,n)$ $=\inf_{\gamma\in{\mathcal{A}}}\mathbb{E}_{\mathbb{Q}}\langle L^{y,\gamma}_{t,s},m_{\mu}(n,z)\rangle\left(\int_{t}^{s}\frac{\gamma^{2}_{u}}{2}\mathrm{d}u\right.$
$\quad\quad+[v(s,Y_{s},z_{s},n_{s})-v(s,y,z,n)]\mathbbm{1}_{\{N_{s}=N_{t}\}}$
[…]
$=:\inf_{\gamma\in{\mathcal{A}}}\left[A^{\gamma}_{1}+A^{\gamma}_{2}+A^{\gamma}_{3}\right].$ (3.1)
Because $v$ is bounded by $\frac{\lambda_{\max}^{2}}{4}$, and because $\mathbb{E}_{\mathbb{Q}}\langle L^{y,\gamma}_{t,s},m_{\mu}(n,z)\rangle\mathbbm{1}_{\{N_{s}>N_{t}\}}\leq\mathbb{P}(N_{s}>N_{t})\leq C(s-t)$, where under $\mathbb{P}$, $(N_{t})$ is a Poisson process of rate $\lambda_{\max}\lVert g\rVert_{\infty}$, we deduce that $|A^{\gamma}_{3}|\leq C(s-t)$ for any control $\gamma\in{\mathcal{A}}$. For the second term, we have using the results of Section 3.1:
$\lvert v(s,Y_{s},z_{s},n_{s})-v(s,y,z,n)\rvert\mathbbm{1}_{\{N_{s}=N_{t}\}}$ $\leq C_{T}\left[\lvert\varphi^{\gamma}_{t,s}(y)-y\rvert+\lvert z_{s}-z\rvert\right]$
$\leq C_{T}(\int_{t}^{s}\lvert\gamma_{u}\rvert\mathrm{d}u+(s-t)\lVert g\rVert_{\infty}).$
Using that $\lvert\gamma_{u}\rvert\leq 1+\gamma^{2}_{u}$, we finally deduce that for any control $\gamma\in{\mathcal{A}}$,
$|A^{\gamma}_{1}|+|A^{\gamma}_{2}|+|A^{\gamma}_{3}|$ $\leq C_{T}(s-t)+C_{T}\mathbb{E}_{\mathbb{Q}}\langle L^{y,\gamma}_{t,s},m_{\mu}(n,z)\rangle\int_{t}^{s}\frac{\gamma^{2}_{u}}{2}\mathrm{d}u.$
Define the following sub-class of admissible control:
${\mathcal{A}}_{0}(t,y,z,n):=\{\gamma\in{\mathcal{A}}:J^{\gamma}(t,y,z,n)\leq J^{0}(t,y,z,n)\}.$
Clearly, we can restrict to this class of control as:
$v(t,y,z,n)=\inf_{\gamma\in{\mathcal{A}}_{0}(t,y,z,n)}J^{\gamma}(t,y,z,n).$
Fix $\epsilon>0$. Let $\gamma\in{\mathcal{A}}_{0}(t,y,z,n)$ be an $\epsilon$-optimal for (3.1). We have:
$\lvert v(t,y,z,n)-v(s,y,z,n)\rvert\leq\epsilon+C_{T}(s-t)+C_{T}\mathbb{E}_{\mathbb{Q}}\langle L^{y,\gamma}_{t,s},m_{\mu}(n,z)\rangle\int_{t}^{s}\frac{\gamma^{2}_{u}}{2}\mathrm{d}u.$
Using the dominated convergence theorem, the left-hand-side is smaller that $3\epsilon$ when $s$ is sufficiently closes to $t$. Altogether, this proves the time continuity of the value function. ∎
3.4 The dynamic programming principle
We now state the general dynamical programming principle which allows from stopping times.
Proposition 3.13.
Grant Assumptions 3.1. Let $\tau$ be a $({\mathcal{F}}^{N}_{t})$ stopping time such that, a.s., $\tau\in[t,T]$. Then
$v(t,y,z,n)=\inf_{\gamma\in{\mathcal{A}}}$ $\mathbb{E}_{\mathbb{Q}}\langle L^{y,\gamma}_{t,\tau},m_{\mu}(n,z)\rangle\left[\int_{t}^{\tau}\frac{\gamma^{2}_{u}}{2}\mathrm{d}u+v(\tau,Y^{t,y,\gamma}_{\tau},Z^{t,y,z,\gamma}_{\tau},n+N_{\tau}-N_{t})\right].$
Proof.
As the value function is continuous (Proposition 3.11), the proof is similar to the proof of Proposition 3.9, with a minor difference: the partition we consider is now a partition of $[0,T]\times\mathbb{R}\times\mathbb{R}_{+}$, to take into account the stopping times. Using Lemma 3.10, the proof is completed as before. ∎
4 Viscosity solution properties and characterization
In this section, to simplify the notations, we write $x=(y,z)\in\mathbb{R}\times\mathbb{R}_{+}$. With this convention, the value function is $v(t,x,n)$, $t\in[0,T]$ and $n\in\mathbb{N}$. By the previous section, we expect that the value function solves formally:
$-\partial_{t}v+H(x,\nabla_{x}v)+\theta_{n}(x)[v-v(t,\psi_{n}(x))]=0,$ (4.1)
with terminal condition
$v(T,x,n)=\kappa\Psi_{\mu}(n,z).$ (4.2)
Here:
$H(x,p):=\frac{1}{2}p_{1}^{2}-B(x)\cdot p,\quad B(x):=\begin{pmatrix}b(y)\\ g(y)\end{pmatrix}\quad\psi_{n}(x):=(0,z,n+1)$
and
$\theta_{n}(x):=\frac{\Phi_{\mu}(n+1,z)}{\Phi_{\mu}(n,z)}g(y).$
We denote by $D:=[0,T]\times\mathbb{R}\times\mathbb{R}_{+}$ and write $BUC(D)$ for the space of all bounded and uniformly continuous functions on $D$. Following [9], we consider the following notion of viscosity solutions.
Definition 4.1.
We say that a function $u:D\times\mathbb{N}\rightarrow\mathbb{R}$ is a viscosity sub-solution of (4.1) if for all $n\in\mathbb{N}$, $(t,x)\mapsto u(t,x,n)\in BUC(D)$ and if for all $(t_{\circ},x_{\circ},n_{\circ})\in D\times\mathbb{N}$, for all $\varphi\in C(D\times\mathbb{N})$ such that $\varphi$ is $C^{1}$ in a open neighborhood of $(t_{\circ},x_{\circ},n_{\circ})$ with $\max(u-\varphi)=(u-\varphi)(t_{\circ},x_{\circ},n_{\circ})=0$, it holds that if $t_{\circ}<T$:
$-\partial_{t}\varphi(t_{\circ},x_{\circ},n_{\circ})+H(x_{\circ},\partial_{x}\varphi(t_{\circ},x_{\circ},n_{\circ}))+\theta_{n_{\circ}}(x_{\circ})\left[u(t_{\circ},x_{\circ},n_{\circ})-u(t_{\circ},\psi_{n_{\circ}}(x_{\circ}))\right]\leq 0.$
We say a function $w:D\times\mathbb{N}\rightarrow\mathbb{R}$ is a viscosity super-solution of (4.1) if for all $n\in\mathbb{N}$, $(t,x)\mapsto w(t,x,n)\in BUC(D)$ and if for all $(t_{\circ},x_{\circ},n_{\circ})\in D\times\mathbb{N}$, for all $\varphi\in C(D\times\mathbb{N})$ such that $\varphi$ is $C^{1}$ in a open neighborhood of $(t_{\circ},x_{\circ},n_{\circ})$ with $\min(w-\varphi)=(w-\varphi)(t_{\circ},x_{\circ},n_{\circ})=0$, it holds that if $t_{\circ}<T$:
$-\partial_{t}\varphi(t_{\circ},x_{\circ},n_{\circ})+H(x_{\circ},\partial_{x}\varphi(t_{\circ},x_{\circ},n_{\circ}))+\theta_{n_{\circ}}(x_{\circ})\left[w(t_{\circ},x_{\circ},n_{\circ})-w(t_{\circ},\psi_{n_{\circ}}(x_{\circ}))\right]\geq 0,$
Finally, we say that $u:D\times\mathbb{R}\rightarrow\mathbb{N}$ is a viscosity solution of (4.1) if it is both a viscosity sub-solution and a viscosity super-solution of this equation.
The main result of this section is
Theorem 4.2.
Under Assumptions 3.1, the value function $v$ is the unique viscosity solution of (4.1) which satisfies the terminal condition (4.2).
We show successively that $v$ is a sub-solution, a super-solution of (4.1) and that a comparison theorem holds. We first need this crucial lemma:
Lemma 4.3.
The process
$\widetilde{N}_{t}:=N_{t}-\int_{0}^{t}\mathbb{E}_{\mathbb{P}}\left[\Lambda\mid\mathcal{F}_{u}^{N}\right]g(Y_{u}^{t,y,\gamma})\mathrm{d}u,\quad t\geq 0$
is an $\mathcal{F}^{N}-$martingale under $\mathbb{P}$.
Proof.
The process is integrable (since $\Lambda\in L^{1}(\mathbb{P})$) and $\mathcal{F}^{N}-$ adapted. Moreover, since
$\mathbb{E}\left[N_{t}-N_{s}\mid\mathcal{F}_{s}^{N}\right]$ $=\mathbb{E}\left[\mathbb{E}\left[N_{t}-N_{s}\mid\sigma(\Lambda,\mathcal{F}_{s}^{N})\right]\mid\mathcal{F}_{s}^{N}\right]=\mathbb{E}\left[\int_{s}^{t}\Lambda g(Y_{u}^{t,y,\gamma})\mathrm{d}u\mid\mathcal{F}_{s}^{N}\right]$
$=\int_{s}^{t}\mathbb{E}\left[\Lambda g(Y_{u}^{t,y,\gamma})\mid\mathcal{F}_{s}^{N}\right]\mathrm{d}u=\int_{s}^{t}\mathbb{E}\left[\mathbb{E}\left(\Lambda g(Y_{u}^{t,y,\gamma})\mid\mathcal{F}_{u}^{N}\right)\mid\mathcal{F}_{s}^{N}\right]\mathrm{d}u$
$=\mathbb{E}\left[\int_{s}^{t}\mathbb{E}\left(\Lambda\mid\mathcal{F}_{u}^{N}\right)g(Y_{u}^{t,y,\gamma})\mathrm{d}u\mid\mathcal{F}_{s}^{N}\right],$
we obtain the result. ∎
4.1 Sub-solution property
In this section, we prove that the value function is a viscosity sub-solution of (4.1). To proceed, we use an equivalent reformulation of the viscosity solutions. We have, following [9]:
Lemma 4.4.
A function $u\in BUC(D)$ for all $n\in\mathbb{N}$ is a viscosity sub-solution of (4.1) if and only if for all $(t_{\circ},x_{\circ},n_{\circ})\in D\times\mathbb{N}$, for all $\varphi\in C(D\times\mathbb{N})$ such that $\varphi$ is $C^{1}$ in an open neighborhood of $(t_{\circ},x_{\circ},n_{\circ})$ and $\max(u-\varphi)=(u-\varphi)(t_{\circ},x_{\circ},n_{\circ})=0$, it holds that if $t_{0}<T$:
$-\partial_{t}\varphi(t_{\circ},x_{\circ},n_{\circ})+H(x_{\circ},\partial_{x}\varphi(t_{\circ},x_{\circ},n_{\circ}))+\theta_{n_{\circ}}(x_{\circ})\left[\varphi(t_{\circ},x_{\circ},n_{\circ})-\varphi(t_{\circ},\psi_{n_{\circ}}(x_{\circ}))\right]\leq 0.$
A similar statement holds for viscosity super-solution.
The only difference with the definition is in the zero order terms, where we have replaced the sub-solution $u$ by the test function $\varphi$. For the proof, see [9, Lem. 2.1]. We now prove that:
Lemma 4.5.
The value function is a viscosity sub-solution of (4.1).
Proof.
With $x_{\circ}:=(y_{\circ},z_{\circ})$, let $(t_{\circ},x_{\circ},n_{\circ})\in D\times\mathbb{N}$ and let $\varphi:D\times\mathbb{N}\rightarrow\mathbb{R}$ a test function as in Lemma 4.4. Let $\gamma\in\mathbb{R}$ be a constant and deterministic control. Let $n_{s}:=n_{\circ}+N_{s}-N_{t_{\circ}}$ and $X_{s}^{t_{\circ},x_{\circ},\gamma}=(Y_{s}^{t_{\circ},y_{\circ},\gamma},Z_{s}^{t_{\circ},y_{\circ},z_{\circ},\gamma}),t_{\circ}\leq s\leq T$. Consider $T_{1}$ the time of the first jump:
$T_{1}=\inf\{s\geq t_{\circ},N_{s}\neq N_{s-}\}.$
As $\gamma$ is a valid control and $s\wedge T_{1}$ a stopping time we have by the dynamical programming principle 3.13:
$\mathbb{E}_{\mathbb{P}}\left[v(t_{\circ},x_{\circ},n_{\circ})-\int_{t_{\circ}}^{s\wedge T_{1}}\frac{\gamma^{2}}{2}\mathrm{d}u-v(s\wedge T_{1},X_{s\wedge T_{1}}^{s_{\circ},x_{\circ},\gamma},n_{s\wedge T_{1}})\right]\leq 0.$
Because $\varphi\geq v$, we deduce that
$\mathbb{E}_{\mathbb{P}}\left[\varphi(t_{\circ},x_{\circ},n_{\circ})-\int_{t_{\circ}}^{s\wedge T_{1}}\frac{\gamma^{2}}{2}\mathrm{d}u-\varphi(s\wedge T_{1},X_{s\wedge T_{1}}^{t_{\circ},x_{\circ},\gamma},n_{s\wedge T_{1}})\right]\leq 0.$
In addition, by Ito’s formula, it holds that:
$\varphi(s\wedge T_{1},X_{s\wedge T_{1}}^{t_{\circ},x_{\circ},\gamma},n_{s\wedge T_{1}}))$ $=\varphi(t_{\circ},x_{\circ},n_{\circ})$
$\!+\int_{t_{\circ}}^{s\wedge T_{1}}[\partial_{t}\varphi+(\partial_{y}\varphi)(b+\gamma)+(\partial_{z}\varphi)g](u,X_{u}^{t_{\circ},x_{\circ},\gamma},n_{u})\mathrm{d}u$
$\!+\int_{t_{\circ}}^{s\wedge T_{1}}\left[\varphi(u,\psi_{n_{u}}(X_{u}^{t_{\circ},x_{\circ},\gamma}))-\varphi(u,X_{u-}^{t_{\circ},x_{\circ},\gamma},n_{u})\right]\theta_{n_{u}}(X_{u}^{t_{\circ},x_{\circ},\gamma})\mathrm{d}u$
$\!+\int_{t_{\circ}}^{s\wedge T_{1}}\left[\varphi(u,\psi_{n_{u}}(X_{u}^{t_{\circ},x_{\circ},\gamma}))-\varphi(u,X_{u-}^{t_{\circ},x_{\circ},\gamma},n_{u})\right]\mathrm{d}\tilde{N}_{u},$
where $\mathrm{d}\tilde{N}_{u}=\mathrm{d}N_{u}-\theta_{n_{u}}(X_{u}^{t_{\circ},x_{\circ},\gamma})\mathrm{d}u=\mathrm{d}N_{u}-\mathbb{E}[\Lambda~{}|~{}{\mathcal{F}}^{N}_{u}]g(Y_{u}^{t_{\circ},y_{\circ},\gamma})\mathrm{d}u$. Taking the expectation with respect to $\mathbb{E}_{\mathbb{P}}$ and using Lemma 4.3, we deduce that
$\frac{1}{s-t_{\circ}}\mathbb{E}_{\mathbb{P}}\left[-\int_{t_{\circ}}^{s\wedge T_{1}}\frac{\gamma^{2}}{2}\mathrm{d}u-\int_{t_{\circ}}^{s\wedge T_{1}}({\mathcal{L}}^{\gamma}\varphi)(u,X_{u}^{t_{\circ},x_{\circ},\gamma},n_{u})\mathrm{d}u\right]\leq 0,$
where the generator ${\mathcal{L}}^{\gamma}$ acting on $\varphi$ is:
$({\mathcal{L}}^{\gamma}\varphi)(t,x,n):=[\partial_{t}\varphi+(\partial_{y}\varphi)(b+\gamma)+(\partial_{z}\varphi)g](t,x,n)+[\varphi(t,\psi_{n}(x))-\varphi(t,x,n)]\theta_{n}(x).$ (4.3)
Note that because $u<T_{1}$, it holds that $n_{u}=n_{\circ}$, and similarly, $X_{u}^{t_{\circ},x_{\circ},\gamma}$ is a deterministic function. We choose $s=t_{\circ}+\frac{1}{k}$, $k\in\mathbb{N}^{*}$ large enough such that $(u,X_{u}^{t_{\circ},x_{\circ},\gamma})$ belongs to the open neighborhood for which the test function $\varphi$ is $C^{1}$, for all $u\leq s\wedge T_{1}$. By applying first the dominated convergence theorem and finally applying the mean value theorem, we deduce that:
$-\frac{\gamma^{2}}{2}-{\mathcal{L}}^{\gamma}\varphi(t_{\circ},x_{\circ},n_{\circ})\leq 0.$
As this is true for all $\gamma$, we choose the optimal value $\gamma=-\partial_{y}\varphi(t_{\circ},x_{\circ},n_{\circ})$ and obtain exactly the required inequality of Lemma 4.4. ∎
4.2 Super-solution property
In this section we prove:
Lemma 4.6.
The value function is a viscosity super-solution of (4.1).
Proof.
We rely similarly on Lemma 4.4. In addition, we note that we can restrict to the class of test functions $\varphi$ for which the minimum is strict, namely $(v-\varphi)(t_{\circ},x_{\circ},n_{\circ})=0$ and:
$(t,x,n)\neq(t_{\circ},x_{\circ},n_{\circ})\implies(v-\varphi)(t,x,n)>0.$ (4.4)
We consider $\varphi$ such a test function. We work towards a contradiction and assume that
$-\partial_{t}\varphi(t_{\circ},x_{\circ},n_{\circ})+H(x_{\circ},\partial_{x}\varphi(t_{\circ},x_{\circ},n_{\circ}))+\theta_{n_{\circ}}(x_{\circ})\left[\varphi(t_{\circ},x_{\circ},n_{\circ})-\varphi(t_{\circ},\psi_{n_{\circ}}(x_{\circ}))\right]<0.$
  1. 1.
    There exists $r>0$ small enough such that
    $-\partial_{t}\varphi+H(x,\partial_{x}\varphi)+\theta_{n}(x)\left[\varphi-\varphi(t,\psi_{n}(x))\right]<0.$ (4.5)
    holds on $B(t_{\circ},x_{\circ},n_{\circ};r)$ which is the ball of center $(t_{\circ},x_{\circ},n_{\circ})$ and of radius $r$. We introduce
    $\widetilde{B}(t_{\circ},x_{\circ},n_{\circ};r):=\left\{(t,\psi_{n}(x))\ \text{for }\ (t,x,n)\in B(t_{\circ},x_{\circ},n_{\circ};r)\right\}\backslash B(t_{\circ},x_{\circ},n_{\circ};r),$
    which is bounded. By (4.4), there exists $\eta>0$ such that
    $\eta:=\min_{(\widetilde{B}\cup\partial B)(t_{\circ},x_{\circ},n_{\circ};r)}(v-\varphi).$ (4.6)
  2. 2.
    Let $\gamma$ be an arbitrary control and $\theta_{\gamma}$ be the first exit time of $B(t_{\circ},x_{\circ},n_{\circ};r)$. By Itô’s formula, we obtain:
    $v(t_{\circ},x_{\circ},n_{\circ})$ $=\varphi(t_{\circ},x_{\circ},n_{\circ})$
    $=\varphi(\theta_{\gamma},X_{\theta_{\gamma}}^{t_{\circ},x_{\circ},\gamma},n_{\theta_{\gamma}})-\int_{t_{\circ}}^{\theta_{\gamma}}{\mathcal{L}}^{\gamma_{u}}\varphi(u,X_{u}^{t_{\circ},x_{\circ},\gamma},n_{u})\mathrm{d}u$
    $\quad-\int_{t_{\circ}}^{\theta_{\gamma}}\left[\varphi(u,\psi_{n_{u}}(X_{u}^{t_{\circ},x_{\circ},\gamma}))-\varphi(u,X_{u}^{t_{\circ},x_{\circ},\gamma},n_{u})\right]\widetilde{N}(\mathrm{d}u),$
    where $\widetilde{N}$ is the martingale process introduced in Lemma 4.3. Note that (4.5) can be rewritten as
    $\sup_{\gamma\in\mathbb{R}}\left[-{\mathcal{L}}^{\gamma}\varphi-\frac{1}{2}\gamma^{2}\right]<0.$
    Therefore:
    $v(t_{\circ},x_{\circ},n_{\circ})$ $\leq\mathbb{E}\left[\varphi\left(\theta_{\gamma},X_{\theta_{\gamma}}^{t_{\circ},x_{\circ},\gamma},n_{\theta_{\gamma}}\right)+\int_{t_{\circ}}^{\theta_{\gamma}}\frac{\gamma^{2}_{u}}{2}\mathrm{d}u\right].$
    Using (4.6), this gives:
    $v(t_{\circ},x_{\circ},n_{\circ})\leq-\eta+\mathbb{E}\left[v\left(\theta_{\gamma},X_{\theta_{\gamma}}^{t_{\circ},x_{\circ},\gamma},n_{\theta_{\gamma}}\right)+\int_{t_{\circ}}^{\theta_{\gamma}}\frac{\gamma_{u}^{2}}{2}\mathrm{d}u\right].$
Since this inequality holds for any control $\gamma$, the latter inequality is in contradiction with Proposition 3.13. ∎
4.3 Comparison theorem
Finally, we provide a comparison principle for the viscosity solutions of (4.1). The main result of this section is:
Proposition 4.7.
Let $u$ be a viscosity sub-solution and let $w$ be a viscosity super-solution of the PDE (4.1), in the sense of Definition 4.1. Assume that $u(T,x,n)\leq w(T,x,n)$ for all $(x,n)\in(\mathbb{R}\times\mathbb{R}_{+})\times\mathbb{N}$. Then
$u(t,x,n)\leq w(t,x,n),\quad\forall t,x,n\in D\times\mathbb{N}.$
In particular, there is at most a bounded and uniformly continuous viscosity solution of (4.1) which satisfies the boundary condition $v(T,x,n)=\kappa\Psi_{n}(x)$.
Proof.
We proceed towards a contradiction and assume that $\theta:=\sup_{D\times\mathbb{N}}u-w>0$.
  1. 1.
    Let $\epsilon,\delta,\beta,\alpha>0$. We consider the auxiliary function
    $\Phi_{n}(t,x,t^{\prime},x^{\prime})$ $=u(t,x,n)-w(t^{\prime},x^{\prime},n)-\frac{1}{2\epsilon}\lVert x-x^{\prime}\rVert^{2}-\frac{1}{2\epsilon}|t-t^{\prime}|^{2}+\beta(t^{\prime}-T)$
    $\quad-\delta F(\lVert x\rVert)-\delta F(\lVert x^{\prime}\rVert)-\alpha n,$
    where $\lVert x\rVert:=\sqrt{y^{2}+z^{2}}$ with $x=(y,z)$ and:
    $F(r):=\log(1+r^{2}),\quad r\geq 0.$
    For all $\epsilon,\delta,\beta,\alpha$ small enough, it holds that
    $\sup_{n\in\mathbb{N}}\sup_{D^{2}}\Phi_{n}>\theta/2.$ (4.7)
    In addition, as $u,w$ are bounded and $\delta>0,\alpha>0$, the $\sup$ is reached for some point $\bar{t},\bar{x},\bar{t}^{\prime},\bar{x}^{\prime},\bar{n}$.
  2. 2.
    From $\Phi_{\bar{n}}(\bar{t},\bar{x},\bar{t}^{\prime},\bar{x}^{\prime})\geq 0$ and $u,w$ bounded, we deduce that for $C=\lVert u\rVert_{\infty}+\lVert w\rVert_{\infty}$:
    $\frac{1}{2\epsilon}\lVert\bar{x}-\bar{x}^{\prime}\rVert^{2}+\frac{1}{2\epsilon}|\bar{t}-\bar{t}^{\prime}|^{2}+\delta F(\lVert\bar{x}\rVert)+\delta F(\lVert\bar{x}^{\prime}\rVert)+\alpha\bar{n}\leq C.$
    In addition, from $\Phi_{\bar{n}}(\bar{t},\bar{x},\bar{t^{\prime}},\bar{x^{\prime}})\geq\Phi_{\bar{n}}(\bar{t^{\prime}},\bar{x^{\prime}},\bar{t^{\prime}},\bar{x^{\prime}})$, we obtain:
    $u(\bar{t},\bar{x},\bar{n})-u(\bar{t^{\prime}},\bar{x^{\prime}},\bar{n})\geq\frac{1}{2\epsilon}\lVert\bar{x}-\bar{x^{\prime}}\rVert^{2}+\frac{1}{2\epsilon}|\bar{t}-\bar{t^{\prime}}|^{2}+\delta F(\lVert\bar{x}\rVert_{2})-\delta F(\lVert\bar{x^{\prime}}\rVert_{2}).$
    For $\rho\geq 0$, let $D_{\rho}=\{(t,x),(t^{\prime},x^{\prime})\in D^{2}:|t-t^{\prime}|^{2}+\lVert x-x^{\prime}\rVert_{2}^{2}\leq\rho\}$, and
    $m^{\alpha}_{u}(\rho)=2\sup\{|u(t,x,n)-u(t^{\prime},x^{\prime},n)|,\quad(t,x),(t^{\prime},x^{\prime})\in D_{\rho},n\in\mathbb{N},n\leq C/\alpha\}.$
    We deduce from $\lVert\bar{x}-\bar{x^{\prime}}\rVert^{2}+|\bar{t}-\bar{t^{\prime}}|^{2}\leq 2C\epsilon$ and from $\lvert\lVert\bar{x}\rVert-\lVert\bar{x^{\prime}}\rVert\rvert\leq\lVert\bar{x}-\bar{x^{\prime}}\rVert$ that:
    $\frac{1}{\epsilon}\lVert\bar{x}-\bar{x^{\prime}}\rVert^{2}+\frac{1}{\epsilon}|\bar{t}-\bar{t^{\prime}}|^{2}\leq m^{\alpha}_{u}(4C\epsilon)+\delta\lVert F^{\prime}\rVert_{\infty}\sqrt{8C\epsilon}.$
    As $(t,x)\mapsto u(t,x,n)$ is uniformly continuous and bounded, the function $m^{\alpha}_{u}:\mathbb{R}_{+}\rightarrow\mathbb{R}_{+}$ is bounded, continuous and $m^{\alpha}_{u}(0+)=m^{\alpha}_{u}(0)=0$. A similar statement holds for $m^{\alpha}_{w}$.
  3. 3.
    Assume that $\bar{t}=T$. We have
    $\Phi_{\bar{n}}(\bar{t},\bar{x},\bar{t^{\prime}},\bar{x^{\prime}})$ $\leq u(\bar{t},\bar{x},\bar{n})-w(\bar{t^{\prime}},\bar{x^{\prime}},\bar{n})$
    $\leq w(\bar{t},\bar{x},\bar{n})-w(\bar{t^{\prime}},\bar{x^{\prime}},\bar{n})$
    $\leq\frac{1}{2}m^{\alpha}_{w}(2C\epsilon).$
  4. 4.
    Similarly, if $\bar{t^{\prime}}=T$, we have:
    $\Phi_{\bar{n}}(\bar{t},\bar{x},\bar{t^{\prime}},\bar{x^{\prime}})\leq\frac{1}{2}m^{\alpha}_{u}(2C\epsilon).$
  5. 5.
    Assume that both $\bar{t}<T$ and $\bar{t^{\prime}}<T$. Consider $\varphi_{u}(t,x):=\frac{1}{2\epsilon}\lVert x-\bar{x^{\prime}}\rVert^{2}+\frac{1}{2\epsilon}|t-\bar{t^{\prime}}|^{2}+\delta F(\lVert x\rVert)$. It holds that $(\bar{t},\bar{x})\in\text{argmax}(u-\varphi_{u})$. So, by the property of viscosity subsolution of $u$, we have:
    $-\frac{\bar{t}-\bar{t^{\prime}}}{\epsilon}+H\left(\bar{x},\frac{\bar{x}-\bar{x^{\prime}}}{\epsilon}+\delta F^{\prime}(\lVert\bar{x}\rVert)\frac{\bar{x}}{\lVert\bar{x}\rVert}\right)+\theta_{\bar{n}}(\bar{x})[u(\bar{t},\bar{x},\bar{n})-u(\bar{t},\psi_{\bar{n}}(\bar{x}))]\leq 0.$
    Similarly, consider $\varphi_{v}(t^{\prime},x^{\prime})=-\frac{1}{2\epsilon}\lVert\bar{x}-x^{\prime}\rVert^{2}-\frac{1}{2\epsilon}|\bar{t}-t^{\prime}|^{2}-\delta F(\lVert x^{\prime}\rVert)+\beta(t^{\prime}-T)$. As $(\bar{t^{\prime}},\bar{x^{\prime}})\in\text{argmin}(w-\varphi_{v})$, we have
    $-\beta-\frac{\bar{t}-\bar{t^{\prime}}}{\epsilon}+H\left(\bar{x^{\prime}},\frac{\bar{x}-\bar{x^{\prime}}}{\epsilon}-\delta F^{\prime}(\lVert\bar{x^{\prime}}\rVert)\frac{\bar{x^{\prime}}}{\lVert\bar{x^{\prime}}\rVert}\right)+\theta_{\bar{n}}(\bar{x^{\prime}})[w(\bar{t^{\prime}},\bar{x^{\prime}},\bar{n})-w(\bar{t^{\prime}},\psi_{\bar{n}}(\bar{x^{\prime}}))]\geq 0.$
    Altogether, combining the two inequalities, we obtain $\beta\leq\Delta_{1}+\Delta_{2}+\Delta_{3}$ where:
    $\Delta_{1}$ $:=H\left(\bar{x^{\prime}},\frac{\bar{x}-\bar{x^{\prime}}}{\epsilon}-\delta F^{\prime}(\lVert\bar{x^{\prime}}\rVert)\frac{\bar{x^{\prime}}}{\lVert\bar{x^{\prime}}\rVert}\right)-H\left(\bar{x},\frac{\bar{x}-\bar{x^{\prime}}}{\epsilon}+\delta F^{\prime}(\lVert\bar{x}\rVert)\frac{\bar{x}}{\lVert\bar{x}\rVert}\right)$
    $\Delta_{2}$ $:=[\theta_{\bar{n}}(\bar{x^{\prime}})-\theta_{\bar{n}}(\bar{x})][w(\bar{t^{\prime}},\bar{x^{\prime}},\bar{n})-w(\bar{t^{\prime}},\psi_{\bar{n}}(\bar{x^{\prime}}))]$
    $\Delta_{3}$ $:=\theta_{\bar{n}}(\bar{x})[w(\bar{t^{\prime}},\bar{x^{\prime}},\bar{n})-w(\bar{t^{\prime}},\psi_{\bar{n}}(\bar{x^{\prime}}))-u(\bar{t},\bar{x},\bar{n})+u(\bar{t},\psi_{\bar{n}}(\bar{x}))].$
    We start with $\Delta_{1}$. We use the explicit shape of $H(x,p)=\frac{1}{2}p_{1}^{2}-B(x)\cdot p$. We have
    $\Delta_{1}$ $=\underbrace{-\frac{\delta}{2}\left[F^{\prime}(\lVert\bar{x}^{\prime}\rVert)\frac{\bar{x}^{\prime}}{\lVert\bar{x}^{\prime}\rVert}+F^{\prime}(\lVert\bar{x}\rVert)\frac{\bar{x}}{\lVert\bar{x}\rVert}\right]_{1}\left[2\frac{\bar{x}-\bar{x}^{\prime}}{\epsilon}+\delta F^{\prime}(\lVert\bar{x}\rVert)\frac{\bar{x}}{\lVert\bar{x}\rVert}-\delta F^{\prime}(\lVert\bar{x}^{\prime}\rVert)\frac{\bar{x}^{\prime}}{\lVert\bar{x}^{\prime}\rVert}\right]_{1}}_{A_{1}}$
    $\quad+\underbrace{\left[B(\bar{x})-B(\bar{x^{\prime}})\right]\cdot\frac{\bar{x}-\bar{x}^{\prime}}{\epsilon}}_{A_{2}}+\underbrace{\delta F^{\prime}(\lVert\bar{x}\rVert)\cdot\frac{B(\bar{x})\cdot\bar{x}}{\lVert\bar{x}\rVert}+\delta F^{\prime}(\lVert\bar{x}^{\prime}\rVert)\frac{B(\bar{x}^{\prime})\cdot\bar{x}^{\prime}}{\lVert\bar{x}^{\prime}\rVert}}_{A_{3}}.$
    We used here the notation $[(y,z)]_{1}=y$. We have:
    $|A_{1}|\leq 2\delta\left(\frac{\sqrt{2C}}{\sqrt{\epsilon}}+\delta\right),\quad|A_{2}|\leq\lVert\nabla B\rVert_{\infty}\left[m^{\alpha}_{u}(4C\epsilon)+\delta\sqrt{8C\epsilon}\right].$
    In addition, there is a constant $\widetilde{C}$ such that $\lvert B(x)\cdot x\rvert\leq\widetilde{C}(\lVert x\rVert^{2}+\lVert x\rVert)$. Using that $\sup_{r\in\mathbb{R}_{+}}\lvert(1+r)F^{\prime}(r)\rvert=1+\sqrt{2}$, we deduce that:
    $|A_{3}|\leq(1+\sqrt{2})\widetilde{C}\delta.$
    Denote by
    $k(\alpha,\epsilon,\delta):=2\delta\left(\frac{\sqrt{2C}}{\sqrt{\epsilon}}+\delta\right)+\lVert\nabla B\rVert_{\infty}\left[m^{\alpha}_{u}(4C\epsilon)+\delta\sqrt{8C\epsilon}\right]+(1+\sqrt{2})\widetilde{C}\delta,$
    we retain that $\Delta_{1}\leq k(\alpha,\epsilon,\delta)$. For $\Delta_{2}$, from Lemma 2.15 and since $g\in C^{1}$, we have:
    $L:=\sup_{n\in\mathbb{N}}\lVert\nabla\theta_{n}\rVert_{\infty}<+\infty.$
    We deduce that:
    $|\Delta_{2}|\leq 2\lVert w\rVert_{\infty}L\sqrt{2C\epsilon}.$
    Finally, we treat $\Delta_{3}$. Let $\bar{x}^{*}=(0,\bar{x}_{2})$ and $\bar{x}^{\prime*}=(0,\bar{x}^{\prime}_{2})$. Using that
    $\Phi_{\bar{n}}(\bar{t},\bar{x},\bar{t}^{\prime},\bar{x}^{\prime})\geq\Phi_{\bar{n}+1}(\bar{t},\bar{x}^{*},\bar{t}^{\prime},\bar{x}^{\prime*}),$
    we deduce that:
    $w(\bar{t^{\prime}},\bar{x^{\prime}},\bar{n})-w(\bar{t^{\prime}},\psi_{\bar{n}}(\bar{x}^{\prime}))-u(\bar{t},\bar{x},\bar{n})+u(\bar{t},\psi_{\bar{n}}(\bar{x}))$ $\leq-\frac{1}{2\epsilon}\lVert\bar{x}-\bar{x}^{\prime}\rVert^{2}+\frac{1}{2\epsilon}|\bar{x}_{2}-\bar{x}^{\prime}_{2}|^{2}$
    $\quad+\delta[F(|\bar{x}_{2}|)-F(\lVert\bar{x}\rVert)]$
    $\quad+\delta[F(|\bar{x}^{\prime}_{2}|)-F(\lVert\bar{x}^{\prime}\rVert)]$
    $\quad+\alpha((n+1)-n).$
    $\leq\alpha.$
    We used crucially that $F$ is non-decreasing and that $|\bar{x}_{2}|=\lVert\bar{x}^{*}\rVert\leq\lVert\bar{x}\rVert$, as just after a jump, the potential is reset to zero.
  6. 6.
    For $\tilde{k}(\alpha,\epsilon,\delta):=k(\alpha,\epsilon,\delta)+2\lVert w\rVert_{\infty}L\sqrt{2C\epsilon}+\alpha$, we have $\beta\leq\tilde{k}(\alpha,\epsilon,\delta)$. As
    $\limsup_{\alpha\downarrow 0}\lim\sup_{\epsilon\downarrow 0}\limsup_{\delta\downarrow 0}\tilde{k}(\epsilon,\delta)=0,$
    there exists $\alpha(\beta)>0$, $\epsilon(\beta,\alpha)>0$ and $\delta(\beta,\alpha,\epsilon)$ such that
    $\forall\alpha\in(0,\alpha(\beta)),\forall\epsilon\in(0,\epsilon(\beta,\alpha)),\forall\delta\in(0,\delta(\beta,\alpha,\epsilon)),\quad\tilde{k}(\alpha,\epsilon,\delta)<\beta.$
    This choice is a contraction with $\beta\leq\tilde{k}(\alpha,\epsilon,\delta)$. Therefore, for those values we necessarily have $\bar{t}=T$ or $\bar{t^{\prime}}=T$. We conclude that
    $\Phi_{\bar{n}}(\bar{t},\bar{x},\bar{t^{\prime}},\bar{x^{\prime}})\leq\frac{1}{2}(m^{\alpha}_{w}+m^{\alpha}_{u})(2C\epsilon).$
    However, when $\epsilon$ is small enough, this contradicts (4.7). This ends the proof.
5 Examples
To illustrate numerically the results, we consider the drift $b(y)=-y$. The controlled process is:
$Y_{t}=y+\int_{0}^{t}(-Y_{u}+\gamma_{u})\mathrm{d}u-\int_{0}^{t}Y_{u-}\mathrm{d}N_{u},\quad t\in[0,T],y\in\mathbb{R}.$
We present two examples using different prior distributions and intensity functions. We first consider an intensity function of the form
$f_{\lambda}(y):=\lambda\exp(2(y-1)),$
and a prior distribution
$\mu:=\sum_{i=1}^{k}p_{i}\delta_{\lambda_{i}},$
where $\sum_{i=1}^{k}p_{i}=1$ and $k\geq 2$. In order to satisfy the assumptions, the function $f_{\lambda}$ can be artificially bounded by a constant $C>0$ so that it does not play an essential role. It is noteworthy that this family of priors is conjugate: the posterior distribution retains the same form, with updated weights $(p_{1},\ldots,p_{k})$. However, within our framework, the dimensionality of the problem is reduced to 2, as opposed to the original $k-1$.
We solve numerically the PDE, using a standard explicit scheme. The analysis of the convergence of this explicit scheme towards the unique viscosity solution of (4.1) can be done using the methodology of [3, Th. 2.1.].
Figure 1: An optimal trajectory with intensity $\lambda\exp(2(y-1))$ and with the prior distribution $\frac{1}{10}\left(\delta_{0}+2\delta_{0.25}+4\delta_{0.5}+2\delta_{0.75}+\delta_{1}\right)$. True value is $\lambda=1$.
In Figure 1, we present a simulated path of the optimal strategy, where the initial measure is given by $m_{0}:=\frac{1}{10}\left(\delta_{0}+2\delta_{0.25}+4\delta_{0.5}+2\delta_{0.75}+\delta_{1}\right)$ and the true parameter value is $\Lambda=1$. The top graphic illustrates the evolution of the potential over time. The second graphic depicts the optimal control, while the third and fourth graphics show the mean and variance of the posterior distribution, respectively. Finally, the bottom-left graphic displays $m_{0}$ as a histogram, and the bottom-right graphic presents $m_{T}$.
Throughout the trajectory, the controller applies a control $t\mapsto\gamma_{t}$ that stays close to one. Near the terminal date, after the final jump, the control becomes small, as increasing the potential further is not worth the cost.
Although the prior is centered at $\lambda=0.5$ with a low probability assigned to $\lambda=1$, the mean of the posterior distribution increases on average. By the end, the posterior distribution assigns the highest probabilities to $\lambda=0.75$ and $\lambda=1$, while excluding $\lambda=0$ and almost completely ruling out $\lambda=0.25$. It is important to note that the posterior probability of $\lambda=0$ immediately becomes zero as soon as a jump is observed.
For the second example, we modify the intensity function as follows:
$f_{\lambda}(y):=\lambda\left[\frac{1}{1+e^{-100(y-1)}}\right].$
This function provides a continuous approximation to the discontinuous function $(\lambda,y)\mapsto\lambda\mathbf{1}_{\{x\geq 1\}}$. In the case of the discontinuous function, for $x<1$, the controller must decide whether to increase the potential to 1 or not. Once the potential reaches 1, it remains fixed thereafter. The continuous version introduced here should exhibit similar overall behavior. For the prior distribution, we choose the uniform probability measure on $[0,2]$:
$\mu:=\mathcal{U}\left([0,2]\right).$
Figure 2: An optimal trajectory with intensity close to $\lambda\mathbf{1}_{\{y\geq 1\}}$ and the prior $\mathcal{U}\left([0,2]\right)$. True value is $\lambda=1$.
In Figure 2, we present a simulated path of the optimal strategy, where the true parameter value is $\Lambda=1$. The graphics illustrate the same functions as in Figure 1, with the distinction that the bottom-left and bottom-right panels display $m_{0}$ and $m_{T}$ as density functions, respectively.
Throughout the trajectory, the controller applies a control $t\mapsto\gamma_{t}$ that increases until the potential reaches a value slightly larger than $Y_{t}=1$. At this point, the control makes a jump to a value close to $\gamma_{t}=1$. Since the drift of the potential is given by $-Y_{t}+\gamma_{t}$, setting $\gamma_{t}=1$ once the potential reaches 1 is optimal, as there is no incentive to exceed this value. As the terminal date approaches, following the final jump, the control value decreases; further increases in potential become unfeasible due to the associated costs.
The prior begins centered on the true value. While the potential remains below one, it remains almost unchanged since we cannot obtain information about $\Lambda$. Once it reaches one, both the mean and variance decrease linearly. When a jump is observed, the mean increases, while the behavior of the variance depends on the specifics of the observation. However, the variance generally decreases over time. Although the prior is symmetric at $\lambda=1$, the observation of a jump excludes the value $\lambda=0$, leading to a density of zero at that point. In contrast, the density at the upper boundary, $\lambda=2$, remains positive, nonetheless close to zero.
Acknowledgments
Nicolas Baradel acknowledges the financial support provided by the Fondation Natixis. Quentin Cormier thanks Carl Graham for useful discussions at different stage of this project, as well as Charles Bertucci and Mete Soner for their help with viscosity solutions.